### v3

# Drinfeld rational fractions for affine Kac–Moody quantum symmetric pairs

Abstract.
We formulate a precise connection between the new Drinfeld presentation of a quantum affine algebra $U_{q}\widehat{{\mathfrak{g}}}$ and the new Drinfeld presentation of affine coideal subalgebras of split type recently discovered by Lu and Wang. In particular, we establish a “factorization formula”, expressing the commuting “Drinfeld–Cartan”-type operators $\Theta_{i,k}$ in a coideal subalgebra in terms of the corresponding Drinfeld generators of $U_{q}\widehat{{\mathfrak{g}}}$, modulo the “Drinfeld positive half” of $U_{q}\widehat{{\mathfrak{g}}}$. We study the spectra of these operators on finite dimensional representations, and describe them in terms of rational functions with an extra symmetry. These results can be seen as the starting point of a $q$-character theory for affine Kac–Moody quantum symmetric pairs. Additionally, we prove a compatibility result linking Lusztig’s and the Lu-Wang-Zhang braid group actions.
Key words and phrases:
Quantum symmetric pairs, coideal subalgebras, $\imath$quantum groups, Kac–Moody algebras, $q$-Onsager algebra, Drinfeld presentation, Drinfeld polynomials, braid group action
2020 Mathematics Subject Classification:
17B37, 17B67, 81R10
The author was supported by the ERC Starting Grant No. 759967 Categorified Donaldson–Thomas Theory, and the EPSRC grant No. EP/W022834/1 Kac–Moody quantum symmetric pairs, KLR algebras and generalized Schur–Weyl duality.
1. Introduction
1.1. A trinity of presentations
It is well known that quantum affine algebras $U_{q}\widehat{{\mathfrak{g}}}$ admit three distinct presentations: the Kac–Moody (also: Drinfeld–Jimbo, Serre) presentation [Jim85, Dri87b], the new Drinfeld (also: current or loop) presentation [Dri87a, Bec94, Dam12, Dam15] and the FRT (also: RTT) presentation [RSTS90]. The first of these can be seen as a quantization of the usual presentation of an affine Kac–Moody Lie algebra, while the second quantizes its construction as a central extension of a loop Lie algebra. Precise isomorphisms between these different presentations where described in [DF93, FM02].
It has recently emerged that the same trinity of presentations exists in the world of quantum symmetric pairs. A (classical) symmetric pair consists of a semisimple Lie algebra, together with a Lie subalgebra fixed under an involution $\theta$. Their classification is equivalent to the classification of real forms of complex semisimple Lie algebras, and can be formulated combinatorially in terms of Satake diagrams. For example, given an appropriate involution, $(\mathfrak{so}_{n},\mathfrak{sl}_{n})$ is a symmetric pair. However, the situation at the quantum level is more nuanced. The usual quantizations of a Lie algebra and its fixed point subalgebra are typically not compatible. The correct quantization of a symmetric pair was proposed by Letzter [Let99, Let02]. According to her definition, a quantum symmetric pair consists of a quantum group and a coideal subalgebra (rather than a Hopf subalgebra), which specializes to $U(\mathfrak{g}^{\theta})$ in the limit $q\mapsto 1$.
The generalization of quantum symmetric pairs to the Kac–Moody case was carried out by Kolb [Kol14]. In particular, he provided a description of coideal subalgebras by generators and Serre-like relations. The RTT presentation of coideal subalgebras was constructed in special cases by Molev, Ragoucy and Sorba [MRS03], and was preceded by earlier work on twisted Yangians [Ols92]. On the other hand, the Drinfeld-type presentation of affine quantum symmetric pairs is a much more recent development. The first steps in this direction were taken by Baseilhac and Kolb in [BK20], where $q$-root vectors and the relations between them were computed in one of the rank $1$ cases (the $q$-Onsager algebra). Their approach was inspired by Damiani’s work for $\widehat{\mathfrak{sl}}_{2}$ [Dam93], and involved the construction of a new braid group action on the $q$-Onsager algebra, using a computer-based proof. A computer-free proof of the existence of this braid group action was given independently by Terwilliger in [Ter18]. A Drinfeld-type presentation for all simply laced split Satake diagrams was later constructed by Lu and Wang [LW21], using a more general braid group action stemming from a Hall algebra realization of coideal subalgebras [LW23]. A generalization to non-simply laced Satake diagrams was obtained by Zhang [Zha22].
1.2. New developments
The theory of affine Kac-Moody quantum symmetric pairs (aKM QSP) is a very dynamic field, with many significant advances in recent years. In order to frame our work in a broader context, we list the most relevant ones below:
  1. i)
    the new Drinfeld presentation for aKM QSP of split type [BK20, LW21, Zha22];
  2. ii)
    relative braid group symmetries on QSP [KP11, WZ23, LWZ23, Zha23], generalizing Lusztig’s braid group action;
  3. iii)
    the Hall algebra approach to aKM QSP [LRW19, LRW23, LW23];
  4. iv)
    universal K-matrices for aKM QSP [AV22a, AV22b];
  5. v)
    generalization of the Kang–Kashiwara–Kim functor, a type of Schur–Weyl duality, to aKM QSP [AP23], forging a link with Khovanov–Lauda–Rouquier-type algebras.
In the present paper we investigate the implications of i) and ii), i.e., the new Drinfeld presentation and braid group action, with view to developing the representation theory of aKM QSP. We expect our work will also exhibit interesting connections to iv) and v) - this is addressed in more detail in §1.6.
1.3. Classical motivations
Based on the seminal works of Bao and Wang [BW18a, BW18b], a general philosophy has emerged that most fundamental algebraic, geometric, and categorical constructions from quantum group theory should be generalizable to quantum symmetric pairs. Further advances in this direction were made in [BK19] and [ES18]. Accordingly, Lu and Wang’s discovery of a Drinfeld-type presentation for aKM QSP can be seen as a major step towards generalizing many classical ideas from the theory of quantum affine algebras. Most importantly, these include:
  1. i)
    the work of Chari and Pressley on Drinfeld polynomials, their relation to spectra of certain commuting operators in the Drinfeld presentation, and their role in the classification of finite dimensional representations [CP91, CP94];
  2. ii)
    the work of Frenkel, Mukhin and Reshetikhin on $q$-characters, associated algorithms, connection to deformed $\mathcal{W}$-algebras and cluster theory [FR99, FM01],
  3. iii)
    Hernandez’s theory of quantum affinizations and the fusion product [Her05, Her07];
  4. iv)
    Nakajima’s geometric construction of finite dimensional representations [Nak01].
The main goal of this paper is to develop a QSP generalization of i), and provide a foundation for ii). A comprehensive study of $q$-characters for quantum symmetric pairs will be presented in a forthcoming paper with Li [LP24]. Some potential implications of our work for iii) are also discussed in §1.6.
1.4. A shift of perspective
In the foundational works of Letzter [Let99, Let02] and Kolb [Kol14], the coideal structure is of primary importance. However, coideal subalgebras admit a description by generators and relations, and may, accordingly, be regarded as an independent algebraic object (often called an $\imath$quantum group), forgetting the embedding into the quantum group. This point of view is more prevalent in recent literature. It is also rather natural for several reasons. Firstly, $\imath$quantum group constructions are not simple restrictions of the corresponding quantum group constructions. For example, the braid group action from [LWZ23] is not a restriction of Lusztig’s braid group action. Secondly, the proofs are often more technically involved and challenging than their quantum group counterparts, exhibiting many features absent from the latter. Examples include a lack of triangular decompositions and canonical raising and lowering operators, the presence of many rank $1$ cases and parameters, as well as non-obvious integral forms.
Nevertheless, from the point of view of representation theory, considering an $\imath$quantum group as embedded inside a quantum group has two key advantages. Firstly, it gives access to restriction functors, which are a natural source of representations. For example, in the case of the $q$-Onsager algebra, essentially every irreducible finite dimensional representation can be obtained as a restriction of an $U_{q}L\mathfrak{sl}_{2}$-representation [IT10]. Secondly, the coproduct endows the category of representations of a coideal subalgebra with a monoidal action. The monoidal structure of the category of finite dimensional representations of a quantum affine algebra has been a subject of intense study, with many important applications to cluster theory [Kas02, Her06, Her10, HL10, HL15]. The corresponding monoidal actions are likely to be equally interesting.
1.5. The main results
These considerations lead to the main problem addressed in this paper:
What is the relation between the Drinfeld presentation of a quantum affine algebra and Lu and Wang’s Drinfeld-type presentation of a coideal subalgebra?
Perhaps rather surprisingly, we establish that these two presentations are, in fact, mutually compatible in a suitable sense. Before stating our results, let us make a few remarks about the basic assumptions and notation. Let ${\mathfrak{g}}$ be a simple Lie algebra of type $\mathsf{ADE}$. We consider the corresponding untwisted quantum affine algebra $U_{q}\widehat{{\mathfrak{g}}}$, and its coideal subalgebras $\mathcal{O}_{q}(\widehat{{\mathfrak{g}}})$ of split type. The latter are also often referred to as generalized $q$-Onsager algebras. The coideal subalgebra $\mathcal{O}_{q}(\widehat{{\mathfrak{g}}})$ depends on two sets of parameters $\mathbf{c}$ and $\mathbf{s}$. The case $\mathbf{s}=0$ is called standard. We use the usual notation for the Drinfeld generators of $U_{q}\widehat{{\mathfrak{g}}}$ (see §8.1 for details). The coideal subalgebra $\mathcal{O}_{q}(\widehat{{\mathfrak{g}}})$, in its Drinfeld presentation, has two main kinds of generators: $A_{i,k}$ (somewhat analogous to $x_{i,k}^{\pm}$ in $U_{q}\widehat{{\mathfrak{g}}}$ and denoted $B_{i,k}$ in [LW21]) and $\Theta_{i,k}$ (analogous to $\phi_{i,k}^{\pm}$ in $U_{q}\widehat{{\mathfrak{g}}}$). When considering the generators $\Theta_{i,k}$ as a series, we use a slightly different normalization, denoted $\boldsymbol{\grave{\Theta}}_{i}(z)$, than in [LW21]. The reason for it is explained in Remark 4.10.
1.5.1. Factorization and coproduct theorems
Of particular importance in the Drinfeld presentation of $U_{q}\widehat{{\mathfrak{g}}}$ is the (almost) commutative subalgebra generated by the elements $\phi_{i,k}^{\pm}$. The spectra of these operators are used to define Drinfeld polynomials and $q$-characters. In a coideal subalgebra $\mathcal{O}_{q}(\widehat{{\mathfrak{g}}})$, there exists an analogous commutative subalgebra generated by elements $\Theta_{i,k}$. Our key result, which we call the factorization theorem, enables us to express the series $\boldsymbol{\grave{\Theta}}_{i}(z)$ in terms of the series $\boldsymbol{\phi}^{\pm}_{i}(z)$, modulo the “Drinfeld positive half” of $U_{q}\widehat{{\mathfrak{g}}}$ (see §8.1 for details).
Theorem 1.1.
Let ${\mathfrak{g}}$ be of type $\mathsf{A}$. Then:
(1.1) $\boldsymbol{\grave{\Theta}}_{i}(z)\equiv\boldsymbol{\phi}_{i}^{-}(z^{-1})\boldsymbol{\phi}_{i}^{+}(Cz){\mod U_{+}}[\negthinspace[z]\negthinspace].$
For non-standard parameters $\mathbf{s}\neq 0$, the statement of Theorem 1.1 is slightly more technical, see Theorem 7.5 for details. The restriction to type $\mathsf{A}$ stems from our use of an explicit combinatorial argument related to braid groups (see §1.5.31.5.4 below). We are confident that this argument can be generalized to other types, which will be explored in future work. Successively more general versions of Theorem 1.1 are proven in Theorems 4.9, 7.5 and Corollary 9.15.
As a consequence of Theorem 1.1, we obtain the following coproduct formula.
Theorem 1.2.
The series $\boldsymbol{\grave{\Theta}}_{i}(z)$ is “approximately group-like”, i.e.,
$\Delta(\boldsymbol{\grave{\Theta}}_{i}(z))\equiv\boldsymbol{\grave{\Theta}}_{i}(z)\otimes\boldsymbol{\grave{\Theta}}_{i}(z)\quad\mod{(U_{-}\otimes U_{+}+U_{+}\otimes U_{-})}[\negthinspace[z]\negthinspace].$
For the proof of Theorem 1.2, see Theorem 7.5 and Corollary 9.15 below.
1.5.2. Drinfeld rational fractions
The Drinfeld presentation of quantum affine algebras is especially well-suited for studying their representation theory. A key role is played by spectra of the series $\boldsymbol{\phi}_{i}^{\pm}(z)$ on finite dimensional representations. By the theorems of Chari–Pressley [CP91, CP94] and Frenkel–Reshetikhin [FR99], they can be described as expansions of rational functions which are ratios of $q^{2}$-shifted polynomials. In the case of the highest weight vector, these polynomials are known as Drinfeld polynomials, and they suffice to classify finite dimensional irreducible representations. To formulate a sensible character theory, the full eigenspace decomposition with respect to the operators $\boldsymbol{\phi}_{i}^{\pm}(z)$ and the corresponding eigenvalues are needed.
As a natural generalization to the world of aKM QSP, we consider the spectra of the series $\boldsymbol{\grave{\Theta}}_{i}(z)$ on finite dimensional representations of $\mathcal{O}_{q}(\widehat{{\mathfrak{g}}})$. Theorem 1.1 enables us to compute these spectra for all representations $V$ obtained via restriction from $U_{q}\widehat{{\mathfrak{g}}}$. More precisely, they can be described in terms of rational functions $F(z)$ with a certain symmetry, which we call Drinfeld rational fractions.
Theorem 1.3.
Consider a generalized eigenspace of $\boldsymbol{\grave{\Theta}}_{i}(z)$ in $V$. The corresponding series of eigenvalues $\boldsymbol{D}(z)$ is the expansion of a rational function of the form
(1.2) $\boldsymbol{D}(z)=\frac{F(q^{-1}z)}{F(qz)},$
where $F(z)$ is a rational function satisfying the $C$-twisted unitarity property:
(1.3) $F(qC^{-1}z^{-1})=F(q^{-1}z)^{-1}.$
Theorem 1.3 (and its close cousins, such as Corollary 5.1 below) is a natural starting point for a $q$-character theory for aKM QSP. For a more precise definition of, as well as a way to compute, Drinfeld rational fractions, see (5.7); for the proof of Theorem 1.3, see Corollary 5.5 below.
1.5.3. Compatibility of braid group actions
It should be emphasized that the braid group action used to construct the Drinfeld presentation for coideal subalgebras in [BK20, LW21] is different from the usual Lusztig braid group action used to construct the Drinfeld presentation for quantum affine algebras. In particular, the former is not a restriction of the latter. Nevertheless, it turns out that, remarkably, the two actions are, roughly speaking, compatible modulo the “Drinfeld positive half” $U_{+}$ of $U_{q}\widehat{{\mathfrak{g}}}$. Not only is this a key ingredient allowing us to generalize Theorem 1.1 from the $q$-Onsager algebra to higher rank coideal subalgebras, but also a very interesting fact on its own. Together with Theorems 1.1 and 1.2, it suggests a general philosophy that although various structures on coideal subalgebras and quantum affine algebras are distinct, they are in fact compatible modulo $U_{+}$.
Let us make the aforementioned compatibility of braid group action more precise. The Drinfeld presentations of both $U_{q}\widehat{{\mathfrak{g}}}$ and $\mathcal{O}_{q}(\widehat{{\mathfrak{g}}})$ rely crucially on braid group operators corresponding to the fundamental weights $\omega_{i}$, applied to the usual Kac–Moody generators. Such generators have the form $B_{i}=F_{i}-c_{i}E_{i}K_{i}^{-1}$ in the coideal case. Let $T_{\omega_{i}}$ denote the usual Lusztig braid group operator, and $\mathbf{T}_{\omega_{i}}$ the corresponding Lu–Wang–Zhang braid group action operator.
Theorem 1.4.
We have
$\mathbf{T}_{\omega_{i}^{\prime}}(B_{i})\equiv T_{\omega_{i}^{\prime}}(B_{i})\mod U_{\neq i,+}.$
Hence the diagram
[…]
commutes modulo $U_{\neq i,+}$.
Here $U_{\neq i,+}$ is the subalgebra spanned by elements of positive $Q_{i}$-degree, where $Q_{i}$ is the root sublattice of $Q$ spanned by all positive roots except $\alpha_{i}$. Theorem 1.4 is proven in §9 (see Corollary 9.7). For the details of the notation, see §8.
1.5.4. A rise in complexity
Our proofs of Theorems 1.11.4 are purely algebraic. However, we would like to point out that their proofs are significantly harder than for some of the corresponding results in the case of quantum affine algebras. For example, the proof of the coproduct formulae for the series $\boldsymbol{\phi}_{i}^{\pm}(z)$ (mod $U_{+}$) relies on relatively straightforward inductions using the Drinfeld presentation. In contrast, the proof for $\boldsymbol{\grave{\Theta}}_{i}(z)$, although also inductive in nature, is much more involved.
The reason for this sharp increase in difficulty can be traced back to the complexity of the Drinfeld presentations themselves. There are two key types of relations which appear both in the original Drinfeld presentation of $U_{q}\widehat{{\mathfrak{g}}}$ and the Lu–Wang presentation of $\mathcal{O}_{q}(\widehat{{\mathfrak{g}}})$. Let us call elements $x_{i,k}^{\pm}$ (resp. $A_{i,k}$) basic, and $h_{i,k},\phi_{i,k}^{\pm}$ (resp. $H_{i,k},\Theta_{i,k}$) diagonal. The first key relation is a commutation relation between basic and diagonal elements, producing further basic elements. In the case of $U_{q}\widehat{{\mathfrak{g}}}$, the RHS of this relation (2.7) involves only one term, while in the relation (2.17) for $\mathcal{O}_{q}(\widehat{{\mathfrak{g}}})$ two terms are present. The second key relation is a commutation relation between basic elements, producing diagonal elements. Again, this relation (2.9) is much easier for $U_{q}\widehat{{\mathfrak{g}}}$, involving only a simple commutator on the LHS and typically one term on the RHS. In contrast, the corresponding relation (2.18) for $\mathcal{O}_{q}(\widehat{{\mathfrak{g}}})$ involves two $q$-commutators on the LHS and typically two terms on the RHS.
1.5.5. Proof outline
The key result, i.e., the factorization theorem (Theorem 1.1) is proven in three stages. The first stage involves rank one, i.e., the $q$-Onsager algebra, and so-called standard embeddings of the coideal subalgebra, i.e., with parameters $\mathbf{s}=0$. The corresponding proof is carried out in §4. The main idea is to decompose the image of the series $\boldsymbol{A}_{+}(z)$ inside $U_{q}\widehat{{\mathfrak{g}}}$ into homogeneous components with respect to the Drinfeld grading. It then suffices to compute the ±$\pm 1$ degree components. Relations (2.17)–(2.18) are used throughout. The second stage of the proof is a generalization to non-standard parameters $\mathbf{s}\neq 0$. Our proof relies on the fact that every non-standard coideal subalgebra can be obtained from a standard one via twisting with a character. Accordingly, it suffices to describe one-dimensional representations of the $q$-Onsager algebra and understand how the coproduct behaves with respect to such representations. This is carried out in §6-7. The final stage involves a generalization to quantum symmetric pairs of higher rank. The key role is played here by rank one subalgebras inside $\mathcal{O}_{q}(\widehat{{\mathfrak{g}}})$ isomorphic to the $q$-Onsager algebra. We study how these subalgebras interact with analogous rank one subalgebras in $U_{q}\widehat{{\mathfrak{g}}}$. Using explicit extended affine Weyl group combinatorics, we establish the braid group action compatibility from Theorem 1.4, which enables us to derive the most general version of the factorization theorem. This is carried out in §8-9.
1.6. Applications and future directions
We will now discuss various applications and insights stemming from our main results.
1.6.1. Affine Cartan subalgebra and representation theory
Our results suggest that the commutative subalgebra generated by the operators $\Theta_{i}$ should be viewed as a kind of “affine Cartan subalgebra” in $\mathcal{O}_{q}(\widehat{{\mathfrak{g}}})$. Moreover, the eigenvalues of these operators are analogous to $\ell$-weights, and the associated eigenspace decomposition provides an analogue of a weight space decomposition.
Nevertheless, the picture is far from clear. Let us mention the most important points. Firstly, since, unlike quantum groups, coideal subalgebras do not possess a triangular decomposition, finding the right replacement for the Cartan subalgebra is quite challenging. This problem was first addressed by Letzter [Let19] for quantum symmetric pairs of finite type. The notion of a Cartan subalgebra proposed therein was used by Wenzl [Wen20] to construct Verma modules for the coideal subalgebra quantizing $U(\mathfrak{so}_{n})$. The previously mentioned lack of canonical raising and lowering operators makes this construction highly non-trivial. Remarkably, the resulting classification of irreducible modules [IK05, Wen20] involves so-called representations of non-classical type, which are not deformations of representations of $U(\mathfrak{so}_{n})$.
It would be interesting to integrate Wenzl’s theory with our approach based on the affine Cartan subalgebra. However, this also appears rather non-trivial. In the case of quantum affine algebras, there is a natural compatibility between the subalgebra generated by the operators $\phi_{i,k}^{\pm}$ from the Drinfeld presentation (we will also refer to it as the “affine Cartan”) and the finite Cartan subalgebra, i.e., $\phi_{i,0}^{\pm}=K_{i}^{\pm 1}$. In contrast, in affine coideal subalgebras, $\Theta_{i,0}$ is merely a scalar and the higher $\Theta_{i,k}$’s do not appear to commute with the generators of Letzter’s proposed Cartan subalgebra.
A related problem, impeding the construction of a sensible highest weight theory for coideal subalgebras, is the lack of a canonical ordering on the eigenvalues of the operators $\Theta_{i,k}$. Finally, we would like to remark that, on restriction representations, the eigenspace decomposition with respect to the affine Cartan subalgebra in $\mathcal{O}_{q}(\widehat{{\mathfrak{g}}})$ is, in general, quite different from (and hence incompatible with) the usual $U_{q}(\widehat{{\mathfrak{g}}})$-weight decomposition with respect to the operators $K_{i}$. Put succinctly: using our results, it is possible to express the eigenvalues of the affine Cartan in $\mathcal{O}_{q}(\widehat{{\mathfrak{g}}})$ in terms of the eigenvalues of the affine Cartan in $U_{q}(\widehat{{\mathfrak{g}}})$, but it is much harder to compute the corresponding eigenvectors.
1.6.2. The symplectic case
Remarkably, the situation in the case of the coideal subalgebras quantizing $U(\mathfrak{sp}_{2n})$ and $U(\widehat{\mathfrak{sp}}_{2n})$ is quite different. Classification results were obtained in [Mol06, GM10], using the RTT presentation. In particular, in the affine case, there exists a classification of irreducibles in terms of Drinfeld polynomials. The nicer behaviour of the symplectic case is, to a large extent, due to the fact that these coideal subalgebras contain some of the Cartan elements $K_{i}$. This facilitates the formulation of a highest weight theory compatible with restriction.
Presently, a Drinfeld-type presentation is only available for aKM QSP of split type. This corresponds to Satake diagrams with no arrows or black dots, and in type $\mathsf{A}$ leads to coideal subalgebras quantizing $U(\widehat{\mathfrak{so}}_{n})$. However, it is natural to expect that a Drinfeld presentation should exist for all aKM QSP, even though the technical complexity of the proofs is likely to increase with the complexity of the Satake diagrams. We also expect that, once the Drinfeld presentation becomes available in new cases, Theorem 1.1 and methods used to establish it will quite straightforwardly generalize to these new cases. In particular, in the symplectic case, corresponding to the Satake diagram with an alternating colouring, the analogue of our theorem should have more direct representation theoretic consequences, allowing us to recover the classification from [GM10] from the point of view of the Drinfeld presentation.
1.6.3. Tridiagonal pairs
The representation theory of coideal subalgebras, outside of the symplectic case, is largely unknown. An exception is the split (i.e. orthogonal) case in rank one, which coincides with the $q$-Onsager algebra. This algebra has been extensively studied by Ito and Terwilliger [IT08, IT10] using completely different combinatorial methods, involving association schemes and tridiagonal pairs. In particular, they formulated the notion of a Drinfeld polynomial for modules over the $q$-Onsager algebra which is, at least on the surface, completely different from ours. Moreover, they obtained an essentially complete classification of irreducible finite dimensional modules.
We believe it will prove fruitful to further explore the interplay between Ito and Terwilliger’s combinatorial approach and our approach based on the Drinfeld presentation. In particular, it is desirable to understand the relation between the different notions of Drinfeld polynomials, and to find an independent classification argument á la Chari and Pressley [CP91]. One may also expect that higher rank quantum symmetric pairs will lead to interesting generalizations of tridiagonal pairs and the associated combinatorics.
1.6.4. Drinfeld coproduct
According to Theorem 1.2, the series $\boldsymbol{\grave{\Theta}}_{i}(z)$ has a group-like property modulo ${U_{+}}[\negthinspace[z]\negthinspace]$. We also prove (Proposition 7.3) that the series $\boldsymbol{A}_{i}{(z)}$ exhibits a kind of “twisted primitive” behaviour. This is completely analogous to the behaviour of the series $\boldsymbol{\phi}_{i}^{\pm}(z)$ and $\boldsymbol{x}_{i}^{\pm}(z)$ in the quantum affine algebra. In the latter case, there exists another coproduct, called the (deformed) Drinfeld coproduct [Her05, Her07], with respect to which such group-like and twisted primitive properties hold exactly. Our results provide further evidence that the same should be true in the world of quantum symmetric pairs. In particular, the theory of quantum affinizations should admit a QSP generalization.
1.6.5. Factorization of the universal K-matrix
Originally, the notion of a $q$-character for quantum affine algebras was formulated in terms of the universal R-matrix [FR99]. Its connection to spectra of the operators $\phi_{i,k}^{\pm}$ was established using the Levendorsky–Soibelman factorization formula [KT94, LSS93, Dam98]. In the QSP case, we are proceeding in the opposite direction, i.e., we propose to define $q$-characters in terms of the spectra of the operators $\Theta_{i,k}$. We expect that an equivalent formulation in terms of the universal K-matrix from [AV22b] is also possible. In particular, there should exist an analogue of the Levendorsky–Soibelman formula for the universal K-matrix. In the finite case, similar factorization results were already established in [WZ23].
1.6.6. Schur–Weyl duality and KLR algebras
It has been known for some time that quantum affine algebras and KLR algebras categorify the same objects. The extra grading on KLR algebras can be seen as explaining the various $q$-deformations appearing in the representation theory of quantum affine algebras, such as $q$-characters and deformations of Grothendieck rings. This link was made precise by Kang, Kashiwara and Kim [KKK18], who constructed “generalized Schur–Weyl duality” functors from categories of modules over KLR algebras to those over quantum affine algebras. Their work was generalized to the QSP case in [AP23]. Accordingly, we expect that $q$-characters for aKM QSP should also have an interesting interpretation in terms of orientifold KLR algebras and the shuffle modules from [Prz23].
1.7. Organization
Sections 2 and 8 consist mainly of recollections of existing material. In Section 3, we establish basic rationality properties of generating series, as well as the important “$C$-symmetry” property. Section 4 is the technical core of the paper, and contains the proof of the first version of the factorization theorem (Theorem 1.1) - the rank one case with parameters $\mathbf{s}=0$. In Section 5, we derive various consequences for spectra of the operators $\Theta_{k}$. In particular, we formulate the notion of Drinfeld rational fractions and prove Theorem 1.3. Sections 6 and 7 extend the factorization theorem to the case of non-standard parameters $\mathbf{s}\neq 0$. We also prove the coproduct theorem (Theorem 1.2). In Section 9, we prove Theorem 1.4 and generalize the factorization theorem to higher ranks.
Acknowledgements
I would like to express my gratitude for all the support I have received. I am particularly indebted to Andrea Appel, who has pointed out to me the importance of rationality properties and one dimensional representations, as well as to Jian-Rong Li for the careful reading of the manuscript and our ongoing collaboration on a related project. I would also like to thank Paul Terwilliger for insightful discussions about the $q$-Onsager algebra, Alexander Molev for correspondence about twisted Yangians, Daniel Orr and Mark Shimozono for providing a reference on the combinatorics of the extended affine Weyl group, as well as Pascal Baseilhac, Weiqiang Wang and Weinan Zhang for comments on the first version of this paper.
2. The $q$-Onsager algebra
Throughout this paper, we assume that $q\in{\mathbb{C}}^{\times}$ is not a root of unity.
2.1. Quantum affine $\mathfrak{sl}_{2}$
The quantum affine algebra $U_{q}\widehat{\mathfrak{sl}}_{2}$ is the associative algebra over $\mathbb{C}$ with generators $e_{i}^{\pm},K_{i}^{\pm 1}$, where $i\in\{0,1\}$, and relations:
$K_{i}K_{i}^{-1}$ $=K_{i}^{-1}K_{i}=1,$
$K_{0}K_{1}$ $=K_{1}K_{0},$
$K_{i}e_{i}^{\pm}$ $=q^{\pm 2}e_{i}^{\pm}K_{i},$
$K_{i}e_{j}^{\pm}$ $=q^{\mp 2}e_{j}^{\pm}K_{i}$ $\quad(i\neq j),$
$[e_{i}^{\pm},e_{j}^{\mp}]$ =±$=\pm\delta_{ij}\frac{K_{i}-K_{i}^{-1}}{q-q^{-1}},$
$\mathsf{Serre}(e_{i}^{\pm},e_{j}^{\pm})$ $=0$ $\quad(i\neq j),$
where $\mathsf{Serre}(x,y)$ denotes the usual $q$-Serre relations, i.e.,
[…]
The algebra $U_{q}\widehat{\mathfrak{sl}}_{2}$ is a Hopf algebra. We will use the coproduct
(2.1) $\Delta(e_{i}^{+})=e_{i}^{+}\otimes 1+K_{i}\otimes e_{i},\quad\Delta(e_{i}^{-})=e_{i}^{-}\otimes K_{i}^{-1}+1\otimes e_{i}^{-},\quad\Delta(K_{i}^{\pm 1})=K_{i}^{\pm 1}\otimes K_{i}^{\pm 1}.$
This choice of coproduct is consistent with [Kol14], and differs from the coproduct in [CP91] by a flip. The counit is given by
$\varepsilon(e_{i}^{\pm})=0,\quad\varepsilon(K_{i}^{\pm 1})=1.$
2.2. The quantum loop algebra
The quantum loop algebra $U_{q}L\mathfrak{sl}_{2}$ is the quotient of $U_{q}\widehat{\mathfrak{sl}}_{2}$ by the relation
(2.2) $K_{1}=K=K_{0}^{-1}.$
Let us recall the Drinfeld presentation of the quantum loop algebra $U_{q}L\mathfrak{sl}_{2}$. By [Dri87a, Bec94], $U_{q}L\mathfrak{sl}_{2}$ is isomorphic to the algebra generated by $x^{\pm}_{k},h_{l},K^{\pm 1}$, where $k\in{\mathbb{Z}},\ l\in{\mathbb{Z}}-\{0\}$, subject to the following relations:
(2.3) $KK^{-1}=$ $\ K^{-1}K=1,$
(2.4) $[h_{k},h_{l}]=$ $\ 0,$
(2.5) $Kh_{k}=$ $\ h_{k}K,$
(2.6) $Kx^{\pm}_{k}=$ $\ q^{\pm 2}x^{\pm}_{k}K,$
(2.7) $[h_{k},x^{\pm}_{l}]=$$\pm\textstyle\frac{[2k]}{k}x^{\pm}_{k+l},$
(2.8) $x^{\pm}_{k+1}x^{\pm}_{l}-q^{\pm 2}x^{\pm}_{l}x^{\pm}_{k+1}=$ $\ q^{\pm 2}x^{\pm}_{k}x^{\pm}_{l+1}-x^{\pm}_{l+1}x^{\pm}_{k},$
(2.9) $[x^{+}_{k},x^{-}_{l}]=$ $\ \textstyle\frac{1}{q-q^{-1}}(\psi_{k+l}-\phi_{k+l}),$
where
(2.10) $\boldsymbol{\psi}(z)=\sum_{k=0}^{\infty}\psi_{k}z^{k}=$ $\ K\exp\left((q-q^{-1})\sum_{k=1}^{\infty}h_{k}z^{k}\right),$
(2.11) $\boldsymbol{\phi}(z)=\sum_{k=0}^{\infty}\phi_{-k}z^{-k}=$ $\ K^{-1}\exp\left(-(q-q^{-1})\sum_{k=1}^{\infty}h_{-k}z^{-k}\right).$
For later convenience, we note that, in particular,
(2.12) $\psi_{0}=K,\quad\phi_{0}=K^{-1},\quad\psi_{1}=(q-q^{-1})Kh_{1},\quad\phi_{-1}=-(q-q^{-1})K^{-1}h_{-1}.$
We also write
(2.13) $\boldsymbol{x}^{+}(z)=\sum_{k=-\infty}^{\infty}x^{+}_{k}z^{k},\quad\boldsymbol{x}^{-}(z)=\sum_{k=-\infty}^{\infty}x^{-}_{k}z^{k}.$
The Serre and Drinfeld presentations are related by the following isomorphism ([Bec94, Proposition 3.7]):
$K\mapsto K,\quad e_{1}^{\pm}\mapsto x^{\pm}_{0},\quad e_{0}^{+}\mapsto-K^{-1}x^{-}_{1},\quad e_{0}^{-}\mapsto-x^{+}_{-1}K.$
We consider $U_{q}L\mathfrak{sl}_{2}$ as a graded algebra with
$\deg x^{\pm}_{k}=\pm 1,\quad\deg h_{k}=0.$
Given $Y\in U_{q}L\mathfrak{sl}_{2}$, let $Y^{(n)}$ denote its homogeneous component of degree $n$. For $i\geq 1$, let $U_{\geq i}$ be the subalgebra of $U_{q}L\mathfrak{sl}_{2}$ spanned by elements of degree at least $i$. If $i=1$, we write $U_{+}=U_{\geq 1}$.
2.3. The $q$-Onsager algebra
The defining relations of the $q$-Onsager algebra, known as the $q$-Dolan–Grady relations, were first considered by Terwilliger [Ter93]. The $q$-Onsager algebra itself was first defined in [Ter05]. In mathematical physics, it was also studied by Baseilhac in [Bas05]. In the framework of [Kol14], the $q$-Onsager algebra is a quantum symmetric pair coideal subalgebra of $U_{q}\widehat{\mathfrak{sl}}_{2}$, corresponding to the Satake diagram of type $A_{1}^{(1)}$ with no black vertices and the trivial automorphism of the underlying affine Dynkin diagram. It is also commonly referred to as the split affine $\imath$quantum group of rank one (e.g., [LW21]).
As in [LW21], we consider a central extension of the $q$-Onsager algebra called the universal $q$-Onsager algebra. It has the following definition by generators and relations.
Definition 2.1.
The universal $q$-Onsager algebra $\mathcal{O}_{q}$ is the algebra generated by $B_{0},B_{1}$ and invertible central elements $\mathfrak{c}_{0},\mathfrak{c}_{1}$ subject to relations:
[…]
It follows from [Kol14, Theorem 8.3] that $Z(\mathcal{O}_{q})={\mathbb{C}}[\mathfrak{c}_{0}^{\pm 1},\mathfrak{c}_{1}^{\pm 1}]$. Given $\mathbf{c}=(c_{0},c_{1})\in({\mathbb{C}}^{\times})^{2}$, we have a central character
(2.15) $\chi_{\mathbf{c}}\colon Z(\mathcal{O}_{q})\to{\mathbb{C}}^{\times},\qquad\mathfrak{c}_{i}\mapsto c_{i}.$
Let $\mathcal{O}_{q}^{\mathbf{c}}$ be the corresponding central reduction, i.e., the quotient of $\mathcal{O}_{q}$ by the ideal generated by $\mathfrak{c}_{i}-c_{i}$. We make no distinction in the notation between elements of $\mathcal{O}_{q}$ and their images in $\mathcal{O}_{q}^{\mathbf{c}}$.
Remark 2.2.
Let us briefly comment on our choice of conventions. In $\mathcal{O}_{q}^{\mathbf{c}}$, the relations (2.14) specialize to the defining relations of [Kol14, Example 7.6]. On the other hand, the dictionary between our conventions and those of [LW21] is given by setting
$v=q,\qquad\mathbb{K}_{i}=q^{2}\mathfrak{c}_{i}.$
2.4. The Lu–Wang presentation
Recently, a Drinfeld-type presentation of $\mathcal{O}_{q}$ was found by Lu and Wang [LW21].
Definition 2.3.
Let ${}^{\mathrm{Dr}}\mathcal{O}_{q}$ be the algebra generated by $H_{m}$ and $A_{r}$, where $m\geq 1$, $r\in{\mathbb{Z}}$, and invertible central elements $\mathfrak{c}_{1},\mathfrak{C}$, subject to the following relations:
(2.16) $[H_{m},H_{n}]$ $=0,$
(2.17) $[H_{m},A_{r}]$ $=\textstyle\frac{[2m]}{m}(A_{r+m}-A_{r-m}\mathfrak{C}^{m}),$
(2.18) $[A_{r},A_{s+1}]_{q^{-2}}-q^{-2}[A_{r+1},A_{s}]_{q^{2}}$ $=\mathfrak{c}_{1}\mathfrak{C}^{r}\Theta_{s-r+1}-q^{-2}\mathfrak{c}_{1}\mathfrak{C}^{r+1}\Theta_{s-r-1}$
$\quad+\mathfrak{c}_{1}\mathfrak{C}^{s}\Theta_{r-s+1}-q^{-2}\mathfrak{c}_{1}\mathfrak{C}^{s+1}\Theta_{r-s-1},$
where $m,n\geq 1$; $r,s\in{\mathbb{Z}}$; and
$1+\sum_{m=1}^{\infty}(q-q^{-1})\Theta_{m}z^{m}=\exp\left((q-q^{-1})\sum_{m=1}^{\infty}H_{m}z^{m}\right).$
In particular, $\Theta_{1}=H_{1}$. By convention, $\Theta_{0}=(q-q^{-1})^{-1}$ and $\Theta_{m}=0$ for $m\leq-1$.
By [LW21, Theorem 2.16], there is an algebra isomorphism
[…]
2.5. Coideal structures
Following [Kol14, Example 7.6], for any $\mathbf{s}=(s_{0},s_{1})\in\mathbb{C}^{2}$, there is an injective algebra homomorphism
(2.20) $\eta_{\mathbf{c},\mathbf{s}}\colon\mathcal{O}_{q}^{\mathbf{c}}\hookrightarrow U_{q}L\mathfrak{sl}_{2},\quad B_{i}\mapsto e_{i}^{-}-c_{i}e_{i}^{+}K_{i}^{-1}+s_{i}K_{i}^{-1},$
giving $\mathcal{O}_{q}^{\mathbf{c}}$ the structure of a right coideal subalgebra of $U_{q}L\mathfrak{sl}_{2}$. Let
$\Delta_{\mathbf{c},\mathbf{s}}=\Delta\circ\eta_{\mathbf{c},\mathbf{s}}\colon\ \mathcal{O}_{q}^{\mathbf{c}}\to\mathcal{O}_{q}^{\mathbf{c}}\otimes U_{q}L\mathfrak{sl}_{2}$
be the associated coproduct. In terms of the Drinfeld presentation, the map (2.20) is given by:
(2.21) $A_{0}\mapsto x^{-}_{0}-c_{1}q^{2}K^{-1}x^{+}_{0}+s_{1}K^{-1},\quad A_{-1}\mapsto-q^{-4}c_{0}^{-1}Kx^{+}_{-1}+x^{-}_{1}+q^{-2}c_{0}^{-1}s_{0}K.$
The choice of a coideal subalgebra structure, or, equivalently, the embedding $\eta_{\mathbf{c},\mathbf{s}}$, gives rise to a right monoidal action
(2.22) $\triangleleft_{\mathbf{c},\mathbf{s}}\colon\operatorname{Mod}^{\operatorname{fd}}_{\mathbf{c}}(\mathcal{O}_{q})\boxtimes\operatorname{Mod^{fd}}(U_{q}L\mathfrak{sl}_{2})\to\operatorname{Mod}^{\operatorname{fd}}_{\mathbf{c}}(\mathcal{O}_{q}),$
where $\operatorname{Mod}^{\operatorname{fd}}_{\mathbf{c}}(\mathcal{O}_{q})$ is the category of finite dimensional $\mathcal{O}_{q}$-modules with a fixed central character (2.15). If $\mathbf{s}=(0,0)$, we call the monoidal action standard.
2.6. One-dimensional modules
Restricting the trivial representation of $U_{q}L\mathfrak{sl}_{2}$ via $\eta_{\mathbf{c},\mathbf{s}}$ gives rise to a continuous family of one-dimensional $\mathcal{O}_{q}^{\mathbf{c}}$-modules $\mathbb{C}_{\mathbf{c},\mathbf{s}}=\eta_{\mathbf{c},\mathbf{s}}^{*}(\mathbb{C})$, parametrised by $\mathbf{s}\in{\mathbb{C}}^{2}$. These modules correspond to the characters
(2.23) $\xi_{\mathbf{c},\mathbf{s}}=\varepsilon\circ\eta_{\mathbf{c},\mathbf{s}}\colon\ \mathcal{O}_{q}^{\mathbf{c}}\ \to\ \mathbb{C},$
explicitly given by
$\xi_{\mathbf{c},\mathbf{s}}(A_{0})=s_{1},\quad\xi_{\mathbf{c},\mathbf{s}}(A_{-1})=q^{-2}c_{0}^{-1}s_{0}.$
Let $\mathcal{O}_{q}^{\mathbf{c},\mathbf{s}}=\eta_{\mathbf{c},\mathbf{s}}(\mathcal{O}_{q})$. By [DK19, Sec. 3.5], the character $\xi_{\mathbf{c},\mathbf{s}}$ induces an algebra isomorphism
$\phi_{\mathbf{c},\mathbf{s}}:\mathcal{O}_{q}^{\mathbf{c},0}\to\mathcal{O}_{q}^{\mathbf{c},\mathbf{s}}$
such that $\eta_{\mathbf{c},\mathbf{s}}=\phi_{\mathbf{c},\mathbf{s}}\circ\eta_{\mathbf{c},0}$. Explicitly, $\phi_{\mathbf{c},\mathbf{s}}$ is given by $\phi_{\mathbf{c},\mathbf{s}}=(\widetilde{\xi}_{\mathbf{c},\mathbf{s}}\otimes\operatorname{id})\circ\Delta|_{\mathcal{O}_{q}^{\mathbf{c},0}}$, where $\widetilde{\xi}_{\mathbf{c},\mathbf{s}}=\xi_{\mathbf{c},\mathbf{s}}\circ(\eta_{\mathbf{c},0}|_{\mathcal{O}_{q}^{\mathbf{c},0}})^{-1}$.
In the following, we will extensively study $\mathcal{O}_{q}$-modules which arise from restricting $U_{q}L\mathfrak{sl}_{2}$-modules via $\eta_{\mathbf{c},\mathbf{s}}$. The following lemma shows that any such module can be realized via a standard action on a one-dimensional representation.
Lemma 2.4.
Let $V\in\operatorname{Mod^{fd}}(U_{q}L\mathfrak{sl}_{2})$. Then, there is an isomorphism of $\mathcal{O}_{q}^{\mathbf{c}}$-modules
$\eta_{\mathbf{c},\mathbf{s}}^{*}(V)\cong\mathbb{C}_{\mathbf{c},\mathbf{s}}\triangleleft_{\mathbf{c},0}V,$
given by the canonical identification of vector spaces $V\cong\mathbb{C}\otimes V$.
Proof.
It is enough to observe that $\eta_{\mathbf{c},\mathbf{s}}=\phi_{\mathbf{c},\mathbf{s}}\circ\eta_{\mathbf{c},0}$, i.e.,
(2.24) $\eta_{\mathbf{c},\mathbf{s}}(Y)=(\widetilde{\xi}_{\mathbf{c},\mathbf{s}}\otimes\operatorname{id})\circ\Delta_{\mathbf{c},0}(Y)$
for any $Y\in\mathcal{O}_{q}^{\mathbf{c}}$. ∎
2.7. $C$-Duality
Let $\operatorname{Mod^{fdw}_{\mathbf{c}}}(\mathcal{O}_{q})$ be the full subcategory of $\mathcal{O}_{q}$-modules $M$ such that:
  1. (a)
    $M$ is finitely generated,
  2. (b)
    $M$ has central character $\chi_{\mathbf{c}}$,
  3. (c)
    there is a joint generalized eigenspace decomposition
    (2.25) $M=\bigoplus_{\bm{\lambda}=(\lambda_{s})_{s\in{\mathbb{Z}}_{\geq 0}}}M_{\bm{\lambda}},\qquad M_{\bf\lambda}=\{v\in M\mid\exists p\ \forall s\in{\mathbb{Z}}_{\geq 0}:(\Theta_{s}-\lambda_{s})^{p}\cdot x=0\},$
    with $M_{\bm{\lambda}}$ finite dimensional.
If $M_{\bm{\lambda}}\neq\{0\}$, we call $\bm{\lambda}$ a weight of $M$. In the context of quantum affine algebras, this is usually referred to as an $\ell$-weight or a pseudo-weight (see, e.g., [Her05, HJ12]).
Lemma 2.5.
There is an algebra anti-automorphism $\tau\colon\mathcal{O}_{q}\to\mathcal{O}_{q}$ sending
$A_{r}\mapsto\mathfrak{C}^{r}A_{-r},\quad H_{m}\mapsto H_{m},\quad\mathfrak{c}_{i}\mapsto\mathfrak{c}_{i},$
inducing a duality
$\operatorname{Mod^{fdw}_{\mathbf{c}}}(\mathcal{O}_{q})\to\operatorname{Mod^{fdw}_{\mathbf{c}}}(\mathcal{O}_{q}),\quad M\mapsto M^{*},$
where $M^{*}$ is the restricted dual with respect to the decomposition (2.25).
Proof.
One directly checks that the defining relations are preserved up to switching the order of multiplication. This holds for (2.17) because
(2.26) $\textstyle\frac{[2m]}{m}\tau(A_{r+m}-A_{r-m}\mathfrak{C}^{m})=$ $\ \textstyle\frac{[2m]}{m}\mathfrak{C}^{r}(A_{-r-m}\mathfrak{C}^{m}-A_{-r+m})$
(2.27) $=$ $\ -[H_{m},\mathfrak{C}^{r}A_{-r}]=\tau([A_{r},H_{m}]).$
To see that (2.18) is also preserved, observe that switching the order of multiplication on the LHS amounts to swapping $r$ and $s$, and the RHS is invariant under such a swap. ∎
We will refer to $M^{*}$ as the $C$-dual of $M$.
3. Rationality
The rationality property of the generating series of quantum loop algebras is due to Beck–Kac [BK96] and Hernandez [Her07] (see also [GTL16]). Below we establish an analogous property for the generating series of the $q$-Onsager algebra. Set
(3.1) $\boldsymbol{A}(z)=\sum_{r=-\infty}^{\infty}A_{r}z^{r},\quad\boldsymbol{A}_{+}(z)=\sum_{r\geq 0}A_{r}z^{r},\quad\boldsymbol{A}_{-}(z)=-\sum_{r\leq-1}A_{r}z^{r}.$
(3.2) $\boldsymbol{\Theta}(z)=\sum_{r\geq 0}\Theta_{r}z^{r},\quad\boldsymbol{\acute{\Theta}}(z)=\frac{1-q^{-2}\mathfrak{C}z^{2}}{1-\mathfrak{C}z^{2}}\boldsymbol{\Theta}(z).$
For convenience, we also abbreviate
$\overline{H}_{1}=[2]^{-1}H_{1},\quad\overline{h}_{\pm 1}=[2]^{-1}h_{\pm 1}.$
Proposition 3.1.
Let $V$ be a module in $\operatorname{Mod^{fdw}_{\mathbf{c}}}(\mathcal{O}_{q})$. Set $C=q^{4}c_{0}c_{1}$. Then, for every weight $\bm{\lambda}$ of $V$:
  1. (1)
    The generating series $\boldsymbol{A}_{\pm}{(z)}\in{\operatorname{End}(V_{\bm{\lambda}})}[\negthinspace[z^{\pm 1}]\negthinspace]$ are the expansions of the same rational function [^1] In the later sections we will typically use the same notation for rational functions and their series expansions where no confusion can arise.
    (3.3) $\mathcal{A}(z)=(1-\operatorname{ad}_{\overline{H}_{1}}z-Cz^{2})^{-1}(A_{0}+CzA_{-1})$
    at $z=0$ and $z=\infty$, respectively.
  2. (2)
    Moreover, the series $\boldsymbol{\acute{\Theta}}(z)\in{\operatorname{End}(V_{\bm{\lambda}})}[\negthinspace[z]\negthinspace]$ is the expansion of the rational function
    (3.4) ${\bm{\vartheta}}(z)=\frac{(1-Cz^{2})\Theta_{0}+c_{1}^{-1}Cz^{2}\left(z^{-1}[A_{-1},\mathcal{A}(z)]_{q^{-2}}-q^{-2}[A_{0},\mathcal{A}(z)]_{q^{2}}\right)}{1-Cz^{2}}$
    at $z=0$.
  3. (3)
    The rational function ${\bm{\vartheta}}(z)$ satisfies the “$C$-symmetry” property:
    (3.5) ${\bm{\vartheta}}(z)={\bm{\vartheta}}(C^{-1}z^{-1}).$
Proof.
(1) Relation (2.17), with $m=1$, implies that
(3.6) $[\overline{H}_{1},\boldsymbol{A}_{\pm}(z)]=$ $\ z^{-1}\boldsymbol{A}_{\pm}(z)-zC\boldsymbol{A}_{\pm}(z)-z^{-1}A_{0}-CA_{-1}.$
Multiplying both sides by $z$ and rearranging the terms, we get
(3.7) $(1-\operatorname{ad}_{\overline{H}_{1}}z-Cz^{2})\boldsymbol{A}_{\pm}(z)=A_{0}+CA_{-1}z.$
(2) Setting $r=-1$ in (2.18), we get, for $s\geq 0$,
(3.8) $[A_{-1},A_{s+1}]_{q^{-2}}-q^{-2}[A_{0},A_{s}]_{q^{2}}=c_{1}(C^{-1}\Theta_{s+2}-q^{-2}\Theta_{s}+\delta_{s=0}\Theta_{0}).$
Also note that, setting $r=s=-1$, relation (2.18) simplifies to
(3.9) $[A_{-1},A_{0}]_{q^{-2}}=q^{-4}c_{0}^{-1}H_{1}=c_{1}C^{-1}\Theta_{1}.$
Hence, summing (3.8) over all $s\geq 0$, we get
(3.10) $-zc_{1}C^{-1}\Theta_{1}\ +$ $\ z^{2}\left(z^{-1}[A_{-1},\boldsymbol{A}_{+}(z)]_{q^{-2}}-q^{-2}[A_{0},\boldsymbol{A}_{+}(z)]_{q^{2}}\right)=$
(3.11) $=$ $\ c_{1}\left((C^{-1}-q^{-2}z^{2})\boldsymbol{\Theta}(z)-C^{-1}\Theta_{1}z+(z^{2}-C^{-1})\Theta_{0}\right).$
This simplifies to
(3.12) $(1-q^{-2}Cz^{2})\boldsymbol{\Theta}(z)=$ $\ (1-Cz^{2})\Theta_{0}$
(3.13) $+z^{2}c_{1}^{-1}C\left(z^{-1}[A_{-1},\boldsymbol{A}_{+}(z)]_{q^{-2}}-q^{-2}[A_{0},\boldsymbol{A}_{+}(z)]_{q^{2}}\right).$
Dividing both sides by $(1-Cz^{2})$, we get (3.4).
(3) On the other hand, for $s\leq-1$ and $r=-1$, relation (2.18) simplifies to
(3.14) $[A_{-1},A_{s+1}]_{q^{-2}}-q^{-2}[A_{0},A_{s}]_{q^{2}}=c_{1}(C^{s}\Theta_{-s}-q^{-2}C^{s+1}\Theta_{-s-2}+C^{-1}(\delta_{s=-1}\Theta_{1}+\delta_{s=-2}\Theta_{0})).$
Summing (3.8) over all $s\leq-1$, we get
(3.15) $z^{-1}c_{1}C^{-1}\Theta_{1}\ +$ $\ z^{-1}[A_{-1},\boldsymbol{A}_{-}(z)]_{q^{-2}}-q^{-2}[A_{0},\boldsymbol{A}_{-}(z)]_{q^{2}}=$
(3.16) $=$ $\ c_{1}\left((1-q^{-2}C^{-1}z^{-2})\boldsymbol{\Theta}(C^{-1}z^{-1})+C^{-1}\Theta_{1}z^{-1}+(z^{-2}C^{-1}-1)\Theta_{0}\right).$
This simplifies to
(3.17) $(1-q^{-2}C^{-1}z^{-2})\boldsymbol{\Theta}(C^{-1}z^{-1})=(1-C^{-1}z^{-2})\Theta_{0}+c_{1}^{-1}\left(z^{-1}[A_{-1},\boldsymbol{A}_{-}(z)]_{q^{-2}}-q^{-2}[A_{0},\boldsymbol{A}_{-}(z)]_{q^{2}}\right).$
Dividing both sides by $(1-C^{-1}z^{-2})$, we get
(3.18) ${\bm{\vartheta}}(C^{-1}z^{-1})=$ $\ \frac{(1-C^{-1}z^{-2})\Theta_{0}+c_{1}^{-1}\left(z^{-1}[A_{-1},\boldsymbol{A}_{-}(z)]_{q^{-2}}-q^{-2}[A_{0},\boldsymbol{A}_{-}(z)]_{q^{2}}\right)}{1-C^{-1}z^{-2}}$
(3.19) $=$ $\ \frac{(1-Cz^{2})\Theta_{0}-c_{1}^{-1}Cz^{2}\left(z^{-1}[A_{-1},\boldsymbol{A}_{-}(z)]_{q^{-2}}-q^{-2}[A_{0},\boldsymbol{A}_{-}(z)]_{q^{2}}\right)}{1-Cz^{2}}.$
Since $\boldsymbol{A}_{+}(z)$ and $-\boldsymbol{A}_{-}(z)$ are both expansions of the same rational function $\mathcal{A}(z)$, the identity ${\bm{\vartheta}}(z)={\bm{\vartheta}}(C^{-1}z^{-1})$ follows. ∎
Remark 3.2.
The $C$-symmetry property does not hold for $\boldsymbol{\Theta}(z)$. This is the first indication that the generators $\acute{\Theta}_{s}$ are more natural from a representation theoretic point of view. We will see further evidence for this later (c.f. Remark 4.10).
4. Factorization of $\boldsymbol{\Theta}(z)$
The goal of this section is to prove our first main result, namely a factorization theorem for $\boldsymbol{\Theta}(z)$ (Theorem 4.9), and interpret it in terms of Drinfeld rational functions (Corollary 5.5). The factorization theorem establishes a connection between the Drinfeld presentation of the $q$-Onsager algebra and the Drinfeld presentation of $U_{q}L\mathfrak{sl}_{2}$. It is key to understanding restricted representations.
More precisely, we are able to express the image of the series $\boldsymbol{\Theta}(z)$ in $U_{q}L\mathfrak{sl}_{2}$ as a product of shifted series $\boldsymbol{\phi}(z)$ and $\boldsymbol{\psi}(z)$, modulo $UX_{+}$. The proof relies on an explicit computation of the homogeneous components of low degree of the images of $\mathcal{O}_{q}$-generators. The main step in the argument is the calculation of $\boldsymbol{A}^{(1)}_{+}(z)$ in §4.3.
For the entirety of §4, we fix a central character (2.15) and the embedding (2.20), assuming that
(4.1) $s_{0}=0=s_{1}.$
This assumption is justified by Lemma 2.4, and will be lifted in §7. For the sake of readability, we make no distinction in the notation between an element of $\mathcal{O}_{q}$ and its image under (2.20). In other words, throughout §4, $A_{i}$ stands for $\eta_{\mathbf{c},0}(A_{i})$, etc.
4.1. Auxiliary relations
We begin by deducing several auxiliary relations in $U_{q}L\mathfrak{sl}_{2}$. By [Nak01, (1.2.7)], the following relation holds:
(4.2) $(z-q^{-2}w)\boldsymbol{\phi}(z)\boldsymbol{x}^{+}(w)=(q^{-2}z-w)\boldsymbol{x}^{+}(w)\boldsymbol{\phi}(z).$
Collecting homogeneous terms, we get
(4.3) $\phi_{-r}x^{+}_{s}+x^{+}_{s-1}\phi_{-r+1}=q^{-2}(x^{+}_{s}\phi_{-r}+\phi_{-r+1}x^{+}_{s-1})$
for $r\geq 1$.
Lemma 4.1.
We have
(4.4) $x^{+}_{k}\phi_{-r}=q^{2}\phi_{-r}x^{+}_{k}+(q^{4}-1)\sum_{s=1}^{r}q^{2(s-1)}\phi_{s-r}x^{+}_{k-s}.$
Proof.
The proof is by induction on $r$. The base case $r=0$ is just the relation (2.6). For $r>0$, we get
(4.5) $x^{+}_{k}\phi_{-r}=$ $\ q^{2}\phi_{-r}x^{+}_{k}-\phi_{-r+1}x^{+}_{k-1}+q^{2}x^{+}_{k-1}\phi_{-r+1}$
(4.6) $=$ $\ q^{2}\phi_{-r}x^{+}_{k}-\phi_{-r+1}x^{+}_{k-1}+q^{2}\left(q^{2}\phi_{-r+1}x^{+}_{k-1}+(q^{4}-1)\sum_{s=1}^{r-1}q^{2(s-1)}\phi_{s-r+1}x^{+}_{k-s-1}\right)$
(4.7) $=$ $\ q^{2}\phi_{-r}x^{+}_{k}+(q^{4}-1)\sum_{s=1}^{r}q^{2(s-1)}\phi_{s-r}x^{+}_{k-s},$
where the first equality follows from (4.3) and the second by induction. ∎
Lemma 4.2.
We have
$x^{+}_{k}\boldsymbol{\phi}(z^{-1})=q^{2}\boldsymbol{\phi}(z^{-1})x^{+}_{k}+(q^{4}-1)z\boldsymbol{\phi}(z^{-1})(1-q^{2}\operatorname{ad}_{\overline{h}_{-1}}z)^{-1}x^{+}_{k-1}.$
Hence
$[x^{+}_{k},\boldsymbol{\phi}(z^{-1})]=\boldsymbol{\phi}(z^{-1})(1-q^{2}\operatorname{ad}_{\overline{h}_{-1}}z)^{-1}(q^{2}-1)(1+\operatorname{ad}_{\overline{h}_{-1}}z)x^{+}_{k}.$
Proof.
By Lemma 4.1, we have
(4.8) $x^{+}_{k}\boldsymbol{\phi}(z^{-1})=$ $\ q^{2}\boldsymbol{\phi}(z^{-1})x^{+}_{k}+(q^{4}-1)\sum_{r\geq 0}\sum_{s=1}^{r}q^{2(s-1)}\phi_{s-r}x^{+}_{k-s}z^{r}$
(4.9) $=$ $\ q^{2}\boldsymbol{\phi}(z^{-1})x^{+}_{k}+(q^{2}-q^{-2})q^{2k}z^{k}\sum_{r\geq 0}\sum_{s=1}^{r}\phi_{s-r}q^{2(s-k)}x^{+}_{k-s}z^{r-k}$
(4.10) $=$ $\ q^{2}\boldsymbol{\phi}(z^{-1})x^{+}_{k}+(q^{2}-q^{-2})q^{2k}z^{k}\boldsymbol{\phi}(z^{-1})\boldsymbol{x}^{+}_{\leq k-1}(q^{-2}z^{-1})$
(4.11) $=$ $\ q^{2}\boldsymbol{\phi}(z^{-1})x^{+}_{k}+(q^{4}-1)z\boldsymbol{\phi}(z^{-1})(1-q^{2}\operatorname{ad}_{\overline{h}_{-1}}z)^{-1}x^{+}_{k-1}.$
Therefore,
(4.12) $[x^{+}_{k},\boldsymbol{\phi}(z^{-1})]=$ $\ (q^{2}-1)\boldsymbol{\phi}(z^{-1})x^{+}_{k}+(q^{4}-1)z\boldsymbol{\phi}(z^{-1})(1-q^{2}\operatorname{ad}_{\overline{h}_{-1}}z)^{-1}x^{+}_{k-1}$
(4.13) $=$ $\ \boldsymbol{\phi}(z^{-1})(1-q^{2}\operatorname{ad}_{\overline{h}_{-1}}z)^{-1}(q^{2}-1)(x^{+}_{k}+x^{+}_{k-1}z)$
(4.14) $=$ $\ \boldsymbol{\phi}(z^{-1})(1-q^{2}\operatorname{ad}_{\overline{h}_{-1}}z)^{-1}(q^{2}-1)(1+\operatorname{ad}_{\overline{h}_{-1}}z)x^{+}_{k},$
as required. ∎
Lemma 4.3.
We have
$z^{-1}[x^{-}_{1},\boldsymbol{\phi}(z^{-1})]_{q^{-2}}-q^{-2}[x^{-}_{0},\boldsymbol{\phi}(z^{-1})]_{q^{2}}=0.$
Proof.
By [Nak01, (1.2.7)],
$(z-q^{-2}w)\boldsymbol{\phi}(z^{-1})\boldsymbol{x}^{-}(w^{-1})=(q^{-2}z-w)\boldsymbol{x}^{-}(w^{-1})\boldsymbol{\phi}(z^{-1}).$
Separating homogeneous terms of degree zero in $w$ and setting $w=1$, we get
$\boldsymbol{\phi}(z^{-1})(zx^{-}_{0}-q^{-2}x^{-}_{1})=q^{-2}(zx^{-}_{0}-x^{-}_{1})\boldsymbol{\phi}(z^{-1}).$
Rearranging the equation above and dividing by $z$ yields the required formula. ∎
4.2. Basic implications of the Lu–Wang relations
We first use the relations (2.16)–(2.18) to express $H_{1}$ in terms of the Drinfeld generators of $U_{q}L\mathfrak{sl}_{2}$.
Lemma 4.4.
We have
(4.15) $H_{1}=$ $\ (Ch_{1}-h_{-1})-q^{-2}c_{1}[x^{+}_{0},x^{+}_{-1}]_{q^{6}}.$
Proof.
Recall from (3.9) that
(4.16) $[A_{-1},A_{0}]_{q^{-2}}=q^{-4}c_{0}^{-1}H_{1}.$
Applying (2.21) to the LHS, the relation becomes
(4.17) $[A_{-1},A_{0}]_{q^{-2}}=$ $\ [x^{-}_{1},x^{-}_{0}]_{q^{-2}}+q^{-2}c_{0}^{-1}c_{1}[Kx^{+}_{-1},K^{-1}x^{+}_{0}]_{q^{-2}}$
(4.18) $\ -q^{2}c_{1}[x^{-}_{1},K^{-1}x^{+}_{0}]_{q^{-2}}-q^{-4}c_{0}^{-1}[Kx^{+}_{-1},x^{-}_{0}]_{q^{-2}}$
(4.19) $=$ $\ -(q^{-6}c_{0}^{-1}c_{1}[x^{+}_{0},x^{+}_{-1}]_{q^{6}}+c_{1}K^{-1}[x^{-}_{1},x^{+}_{0}]+q^{-4}c_{0}^{-1}K[x^{+}_{-1},x^{-}_{0}])$
(4.20) $=$ $\ -q^{-6}c_{0}^{-1}c_{1}[x^{+}_{0},x^{+}_{-1}]_{q^{6}}+c_{1}h_{1}-q^{-4}c_{0}^{-1}h_{-1}.$
Note that the first term on the RHS of (4.17) vanishes thanks to (2.8). ∎
Next, we compute $A_{r}^{(-1)}$. Relation (2.17) (with $m=1$) implies that
(4.21) $A_{r}=[\overline{H}_{1},A_{r-1}]+CA_{r-2}.$
Lemma 4.5.
We have:
[…]
(4.23) $A_{r}^{(-1)}=$ $\ x^{-}_{-r}$ $(r\in{\mathbb{Z}}).$
Proof.
Recall from (2.21) that, assuming (4.1), generators $A_{-1}$ and $A_{0}$ only have terms of degree ±$\pm 1$. The first equality (4.22) then follows from (4.21), by downward and upward inductions, using the fact that $H_{1}$ has only terms of degree 00 and $2$ (see (4.15)).
The second equality holds for $r=-1,0$ by (2.21). By induction and Lemma 4.15, we get
(4.24) $A_{r}^{(-1)}=$ $\ [\overline{H}_{1}^{(0)},A_{r-1}^{(-1)}]+CA_{r-2}^{(-1)}$
(4.25) $=$ $\ [C\overline{h}_{1}-\overline{h}_{-1}),x^{-}_{-r+1}]+Cx^{-}_{-r+2}$
(4.26) $=$ $\ -Cx^{-}_{-r+2})+x^{-}_{-r}+Cx^{-}_{-r+2}$
(4.27) $=$ $\ x^{-}_{-r}$
for $r\geq 1$. A similar argument using the relation $A_{r}=C^{-1}(A_{r+2}-[\overline{H}_{1},A_{r+1}])$ and downward induction establishes (4.23) for $r\leq-2$. ∎
In terms of generating series, (4.23) can be rephrased as
(4.28) $\boldsymbol{A}^{(-1)}(z)=\boldsymbol{x}^{-}(z^{-1}).$
4.3. $\boldsymbol{A}^{(1)}_{+}(z)$ calculation
Abbreviate
(4.29) $\Omega=(1+\operatorname{ad}_{\overline{h}_{-1}}z)(1-C\operatorname{ad}_{\overline{h}_{1}}z).$
Lemma 4.6.
We have
(4.30) $\boldsymbol{A}^{(1)}_{+}(z)=$ $\ -c_{1}\boldsymbol{\phi}(z^{-1})\Omega^{-1}\big{(}(q^{2}-q^{-2})\boldsymbol{x}^{+}_{\leq-2}(q^{-2}z^{-1})+q^{2}x^{+}_{0}+q^{4}x^{+}_{-1}z\big{)}.$
Proof.
First, observe that
(4.31) $\left((1-\operatorname{ad}_{\overline{H}_{1}}z-Cz^{2})\boldsymbol{A}_{+}(z)\right)^{(1)}=$ $\ (1-\operatorname{ad}_{\overline{H}_{1}^{(0)}}z-Cz^{2})\boldsymbol{A}^{(1)}_{+}(z)-z\operatorname{ad}_{\overline{H}_{1}^{(2)}}\boldsymbol{A}^{(-1)}_{+}(z)$
(4.32) $=$ $\ \Omega\boldsymbol{A}^{(1)}_{+}(z)-z\operatorname{ad}_{\overline{H}_{1}^{(2)}}\boldsymbol{A}^{(-1)}_{+}(z),$
(4.33) $(A_{0}+CA_{-1}z)^{(1)}=$ $\ -c_{1}(Kx^{+}_{-1}z+q^{2}K^{-1}x^{+}_{0}).$
Hence, by (3.7), we get
(4.34) $\Omega\boldsymbol{A}^{(1)}_{+}(z)-z\operatorname{ad}_{\overline{H}_{1}^{(2)}}\boldsymbol{A}^{(-1)}_{+}(z)=-c_{1}(Kx^{+}_{-1}z+q^{2}K^{-1}x^{+}_{0}).$
We now calculate the second term on the LHS. We have
(4.35) $\operatorname{ad}_{\overline{H}_{1}^{(2)}}\boldsymbol{A}^{(-1)}_{+}(z)=$ $\ -q^{-2}\frac{c_{1}}{[2]}\sum_{r\geq 0}\big{[}[x^{+}_{0},x^{+}_{-1}]_{q^{6}},x^{-}_{-r}\big{]}z^{r}.$
Let us compute the terms in the sum on the RHS. We get
(4.36) $x^{+}_{0}x^{+}_{-1}x^{-}_{-r}=$ $\ x^{+}_{0}(x^{-}_{-r}x^{+}_{-1}-(q-q^{-1})^{-1}\phi_{-r-1})$
(4.37) $=$ $\ -(q-q^{-1})^{-1}((\phi_{-r}-\delta_{r=0}\psi_{0})x^{+}_{-1}+x^{+}_{0}\phi_{-r-1})+x^{-}_{-r}x^{+}_{0}x^{+}_{-1},$
(4.38) $x^{+}_{-1}x^{+}_{0}x^{-}_{-r}=$ $\ x^{+}_{-1}(x^{-}_{-r}x^{+}_{0}-(q-q^{-1})^{-1}(\phi_{-r}-\delta_{r=0}\psi_{0}))$
(4.39) $=$ $\ -(q-q^{-1})^{-1}(\phi_{-r-1}x^{+}_{0}+x^{+}_{-1}(\phi_{-r}-\delta_{r=0}\psi_{0}))+x^{-}_{-r}x^{+}_{-1}x^{+}_{0}$
(4.40) $=$ $\ -q^{-2}(q-q^{-1})^{-1}((\phi_{-r}-\delta_{r=0}\psi_{0})x^{+}_{-1}+x^{+}_{0}\phi_{-r-1})+x^{-}_{-r}x^{+}_{-1}x^{+}_{0},$
where the last equality follows from (4.3). Hence, using Lemma 4.1, we get
(4.41) $\big{[}[x^{+}_{0},x^{+}_{-1}]_{q^{6}},x^{-}_{-r}\big{]}=$ $\ q^{2}[2](\phi_{-r}x^{+}_{-1}+x^{+}_{0}\phi_{-r-1}-\delta_{r=0}Kx^{+}_{-1})$
(4.42) $=$ $\ [2]\big{(}q^{4}\phi_{-r-1}x^{+}_{0}+q^{6}\phi_{-r}x^{+}_{-1}+(q^{4}-1)\sum_{s=2}^{r+1}q^{2s}\phi_{s-r-1}x^{+}_{-s}-\delta_{r=0}q^{2}Kx^{+}_{-1}\big{)}.$
Combining the equality above with (4.35), we obtain
(4.43) $\operatorname{ad}_{\overline{H}_{1}^{(2)}}\boldsymbol{A}^{(-1)}_{+}(z)=$ $\ c_{1}Kx^{+}_{-1}-q^{-2}c_{1}\sum_{r\geq 0}(q^{4}\phi_{-r-1}x^{+}_{0}+q^{6}\phi_{-r}x^{+}_{-1})z^{r}$
(4.44) $\ -c_{1}(q^{2}-q^{-2})\sum_{r\geq 0}\sum_{s=2}^{r+1}q^{2s}\phi_{s-r-1}x^{+}_{-s}z^{r}.$
(4.45) $=$ $\ c_{1}Kx^{+}_{-1}+c_{1}q^{2}K^{-1}x^{+}_{0}z^{-1}-c_{1}q^{2}\boldsymbol{\phi}(z^{-1})x^{+}_{0}z^{-1}$
(4.46) $\ -c_{1}q^{4}\boldsymbol{\phi}(z^{-1})x^{+}_{-1}-c_{1}(q^{2}-q^{-2})\boldsymbol{\phi}(z^{-1})\boldsymbol{x}^{+}_{\leq-2}(q^{-2}z^{-1})z^{-1}.$
Finally, combining the equality above with (4.34), we get
(4.47) $\Omega\boldsymbol{A}^{(1)}_{+}(z)=$ $\ -c_{1}\boldsymbol{\phi}(z^{-1})\big{(}(q^{2}-q^{-2})\boldsymbol{x}^{+}_{\leq-2}(q^{-2}z^{-1})+q^{2}x^{+}_{0}+q^{4}x^{+}_{-1}z\big{)}.$
Dividing both sides by $\Omega$ and using the fact that $\Omega$ commutes with $\boldsymbol{\phi}(z^{-1})$ completes the proof. ∎
In light of Lemma 4.6, we define
(4.48) $\mathbb{A}=-c_{1}\Omega^{-1}\big{(}(q^{2}-q^{-2})\boldsymbol{x}^{+}_{\leq-2}(q^{-2}z^{-1})+q^{2}x^{+}_{0}+q^{4}x^{+}_{-1}z\big{)},$
so that $\boldsymbol{A}^{(1)}_{+}(z)=\boldsymbol{\phi}(z^{-1})\mathbb{A}$.
4.4. Proof of the factorization theorem
Throughout this subsection the symbol $\equiv$” denotes equalities holding modulo $U_{+}$.
Lemma 4.7.
We have
(4.49) $(1-q^{-2}Cz^{2})\boldsymbol{\Theta}(z)\equiv$ $\ \frac{(K-CK^{-1}z^{2})\boldsymbol{\phi}(z^{-1})}{q-q^{-1}}$
(4.50) $\ -q^{2}c_{0}z\boldsymbol{\phi}(z^{-1})[\mathbb{A},(1+q^{2}\operatorname{ad}_{\overline{h}_{-1}}z)x_{1}^{-}]\quad\mod{U_{+}}[\negthinspace[z]\negthinspace].$
Proof.
Recall from (3.12) that
$(1-q^{-2}Cz^{2})\boldsymbol{\Theta}(z)=(1-Cz^{2})\Theta_{0}+z^{2}c_{1}^{-1}C\left(z^{-1}[A_{-1},\boldsymbol{A}_{+}(z)]_{q^{-2}}-q^{-2}[A_{0},\boldsymbol{A}_{+}(z)]_{q^{2}}\right).$
The elements $A_{-1}$ and $A_{0}$ only contain terms of degree ±$\pm 1$, whereas $\boldsymbol{A}_{+}(z)$ has additional terms of higher degree. Since we are working modulo $U_{+}$, we can discard all the homogeneous components $\boldsymbol{A}^{(l)}_{+}(z)$ with $l\geq 2$. Therefore,
(4.51) $z^{-1}[A_{-1},\boldsymbol{A}_{+}(z)]_{q^{-2}}-q^{-2}[A_{0},\boldsymbol{A}_{+}(z)]_{q^{2}}\equiv$ $\ (z^{-1}A_{-1}-q^{-2}A_{0})(\boldsymbol{A}^{(-1)}_{+}(z)+\boldsymbol{A}^{(1)}_{+}(z))$
(4.52) $\ +(\boldsymbol{A}^{(-1)}_{+}(z)+\boldsymbol{A}^{(1)}_{+}(z))(-q^{-2}z^{-1}A_{-1}+A_{0}).$
Clearly, we can ignore terms of degree $2$ above. Next, we show that terms of degree $-2$ cancel out, i.e.,
(4.53) $(z^{-1}A_{-1}^{(-1)}-q^{-2}A_{0}^{(-1)})\boldsymbol{A}^{(-1)}_{+}(z)+\boldsymbol{A}^{(-1)}_{+}(z)(-q^{-2}z^{-1}A_{-1}^{(-1)}+A_{0}^{(-1)})=0.$
Indeed, computing the LHS of (4.53), we get
(4.54) $\mbox{LHS}=(z^{-1}x^{-}_{1}-q^{-2}x^{-}_{0})\boldsymbol{x}^{-}_{\leq 0}(z^{-1})+\boldsymbol{x}^{-}_{\leq 0}(z^{-1})(q^{-2}z^{-1}x^{-}_{1}+x^{-}_{0}).$
The $z^{s}$-term in the resulting series is equal to
$(x^{-}_{1}x^{-}_{-s-1}-q^{-2}x^{-}_{0}x^{-}_{-s})+(-q^{-2}x^{-}_{-s-1}x^{-}_{1}+x^{-}_{-s}x^{-}_{0})$
if $s\geq 0$ and $[x^{-}_{1},x^{-}_{0}]_{q^{-2}}$ if $s=-1$. In both cases, the expression vanishes thanks to relation (2.8), with $k=0$ and $l=-s-1$.
It follows that
(4.55) $\ \ z^{-1}[A_{-1},\boldsymbol{A}_{+}(z)]_{q^{-2}}-q^{-2}[A_{0},\boldsymbol{A}_{+}(z)]_{q^{2}}\equiv$ $\ (z^{-1}A_{-1}^{(1)}-q^{-2}A_{0}^{(1)})\boldsymbol{A}^{(-1)}_{+}(z)$
(4.56) $\ +(z^{-1}A_{-1}^{(-1)}-q^{-2}A_{0}^{(-1)})\boldsymbol{A}^{(1)}_{+}(z)$
(4.57) $\ +\boldsymbol{A}^{(-1)}_{+}(z)(-q^{-2}z^{-1}A_{-1}^{(1)}+A_{0}^{(1)})$
(4.58) $\ +\boldsymbol{A}^{(1)}_{+}(z)(-q^{-2}z^{-1}A_{-1}^{(-1)}+A_{0}^{(-1)}).$
Our next step is to compute the four summands on the RHS above.
For $s\geq 0$, we have
(4.59) $(A_{-1}^{(1)}x^{-}_{-s-1}-q^{-2}A_{0}^{(1)}x^{-}_{-s})=$ $\ (-q^{-4}c_{0}^{-1}Kx^{+}_{-1}x^{-}_{-s-1}+c_{1}K^{-1}x^{+}_{0}x^{-}_{-s}),$
(4.60) $(-q^{-2}x^{-}_{-s-1}A_{-1}^{(1)}+x^{-}_{-s}A_{0}^{(1)})=$ $\ (q^{-4}c_{0}^{-1}Kx^{-}_{-s-1}x^{+}_{-1}-c_{1}K^{-1}x^{-}_{-s}x^{+}_{0}),$
and summing the two terms above yields
$\frac{K\phi_{-s-2}-CK^{-1}(\phi_{-s}-\delta_{s=0}\psi_{0})}{c_{1}^{-1}C(q-q^{-1})}.$
We also have
$A_{-1}^{(1)}x^{-}_{0}-q^{-2}x^{-}_{0}A_{-1}^{(1)}=-q^{-4}c_{0}^{-1}K[x^{+}_{-1},x^{-}_{0}]=(q-q^{-1})^{-1}c_{1}C^{-1}K\phi_{-1}.$
Putting the equalities above into a series, we get
(4.62) $(z^{-1}A_{-1}^{(1)}-q^{-2}A_{0}^{(1)})\boldsymbol{A}^{(-1)}_{+}(z)\ +$ $\ \boldsymbol{A}^{(-1)}_{+}(z)(-q^{-2}z^{-1}A_{-1}^{(1)}+A_{0}^{(1)})\equiv$
(4.63) $\equiv$ $\ \frac{(K-CK^{-1}z^{2})\boldsymbol{\phi}(z^{-1})+(Cz^{2}-1)}{c_{1}^{-1}Cz^{2}(q-q^{-1})}.$
Moreover, by Lemma 4.6, we have
(4.64) $(z^{-1}A_{-1}^{(-1)}-q^{-2}A_{0}^{(-1)})\boldsymbol{A}^{(1)}_{+}(z)=$ $\ (z^{-1}A_{-1}^{(-1)}-q^{-2}A_{0}^{(-1)})\boldsymbol{\phi}(z^{-1})\mathbb{A}$
(4.65) $=$ $\ (x_{1}^{-}z^{-1}-q^{-2}x_{0}^{-})\boldsymbol{\phi}(z^{-1})\mathbb{A}$
(4.66) $=$ $\ \boldsymbol{\phi}(z^{-1})(q^{-2}x_{1}^{-}z^{-1}-x_{0}^{-})\mathbb{A},$
where $\mathbb{A}$ is as in (4.48), and the third equality follows from Lemma 4.3. On the other hand,
$\boldsymbol{A}^{(1)}_{+}(z)(-q^{-2}z^{-1}A_{-1}^{(-1)}+A_{0}^{(-1)})=\boldsymbol{\phi}(z^{-1})\mathbb{A}(-q^{-2}x^{-}_{1}z^{-1}+x^{-}_{0}).$
Hence
(4.67) $(z^{-1}A_{-1}^{(-1)}-q^{-2}A_{0}^{(-1)})\boldsymbol{A}^{(1)}_{+}(z)\ +$ $\ \boldsymbol{A}^{(1)}_{+}(z)(-q^{-2}z^{-1}A_{-1}^{(-1)}+A_{0}^{(-1)})=$
(4.68) $=$ $\ -q^{-2}z^{-1}\boldsymbol{\phi}(z^{-1})[\mathbb{A},(1+q^{2}\operatorname{ad}_{\overline{h}_{-1}}z)x_{1}^{-}].$
Substituting (4.62) and (4.67) back into (4.55) and (3.12) completes the proof. ∎
Lemma 4.8.
We have
(4.69) $(1-q^{2}\operatorname{ad}_{\overline{h}_{-1}}z)\mathbb{A}=-c_{1}q^{2}(1-\operatorname{ad}_{\overline{h}_{-1}}z)\boldsymbol{x}^{+}_{\geq 0}(Cz)$
Proof.
We begin with an auxiliary calculation. Abbreviate
(4.70) $M=$ $\ -[(q^{2}-q^{-2})\boldsymbol{x}^{+}_{\leq-2}(q^{-2}z^{-1})+q^{2}x^{+}_{0}+q^{4}x^{+}_{-1}z].$
We have
(4.71) $-(1-q^{2}\operatorname{ad}_{\overline{h}_{-1}}z)(q^{2}-q^{-2})\boldsymbol{x}^{+}_{\leq-2}(q^{-2}z^{-1})=$ $\ (q^{2}-q^{6})x^{+}_{-2}z^{2}$
(4.72) $-(1-q^{2}\operatorname{ad}_{\overline{h}_{-1}}z)(q^{2}x^{+}_{0}+q^{4}x^{+}_{-1}z)=$ $\ q^{6}x^{+}_{-2}z^{2}-q^{2}x^{+}_{0}.$
Hence
(4.73) $(1-q^{2}\operatorname{ad}_{\overline{h}_{-1}}z)M=q^{2}x^{+}_{-2}z^{2}-q^{2}x^{+}_{0}=(1-\operatorname{ad}_{\overline{h}_{-1}}^{2}z^{2})(-q^{2}x^{+}_{0})$
and so
(4.74) $(1-q^{2}\operatorname{ad}_{\overline{h}_{-1}}z)\Omega^{-1}M=-q^{2}\frac{1-\operatorname{ad}_{\overline{h}_{-1}}z}{1-C\operatorname{ad}_{\overline{h}_{1}}z}x^{+}_{0}.$
Recall that, according to Lemma 4.6,
(4.75) $\mathbb{A}=c_{1}\Omega^{-1}M.$
Therefore,
(4.76) $(1-q^{2}\operatorname{ad}_{\overline{h}_{-1}}z)\mathbb{A}=$ $\ c_{1}(1-q^{2}\operatorname{ad}_{\overline{h}_{-1}}z)\Omega^{-1}M$
(4.77) $=$ $\ -q^{2}c_{1}\frac{1-\operatorname{ad}_{\overline{h}_{-1}}z}{1-C\operatorname{ad}_{\overline{h}_{1}}z}x^{+}_{0}$
(4.78) $=$ $\ -c_{1}q^{2}(1-\operatorname{ad}_{\overline{h}_{-1}}z)\boldsymbol{x}^{+}_{\geq 0}(Cz),$
yielding the desired formula. The first equality follows from (4.75), the second from (4.74) and the third from (2.7). ∎
We now have all the ingredients to prove the main result of this section.
Theorem 4.9 (Factorization theorem).
Assuming that $s_{0}=s_{1}=0$, we have
(4.79) $\boldsymbol{\Theta}(z)\equiv\frac{1-Cz^{2}}{(q-q^{-1})(1-q^{-2}Cz^{2})}\boldsymbol{\phi}(z^{-1})\boldsymbol{\psi}(Cz)\mod{U_{+}}[\negthinspace[z]\negthinspace].$
Proof.
We have the following chain of equalities:
(4.80) $-q^{2}c_{0}z\boldsymbol{\phi}(z^{-1})[\mathbb{A},(1\ +$ $\ q^{2}\operatorname{ad}_{\overline{h}_{-1}}z)x_{1}^{-}]=-q^{2}c_{0}z\boldsymbol{\phi}(z^{-1})[(1-q^{2}\operatorname{ad}_{\overline{h}_{-1}}z)\mathbb{A},x_{1}^{-}]$
(4.81) $=$ $\ Cz\boldsymbol{\phi}(z^{-1})[(1-\operatorname{ad}_{\overline{h}_{-1}}z)\boldsymbol{x}^{+}_{\geq 0}(Cz),x^{-}_{1}]$
(4.82) $=$ $\ Cz\boldsymbol{\phi}(z^{-1})[\boldsymbol{x}^{+}_{\geq 0}(Cz),(1+\operatorname{ad}_{\overline{h}_{-1}}z)x^{-}_{1}]$
(4.83) $=$ $\ \frac{Cz\boldsymbol{\phi}(z^{-1})}{q-q^{-1}}\big{(}(Cz)^{-1}\boldsymbol{\psi}_{\geq 1}(Cz)-z\boldsymbol{\psi}(Cz)+zK^{-1}\big{)}$
(4.84) $=$ $\ \frac{\boldsymbol{\phi}(z^{-1})}{q-q^{-1}}\big{(}(1-Cz^{2})\boldsymbol{\psi}(Cz)-K+Cz^{2}K^{-1}\big{)}.$
The first and third equalities follow from the fact that $\operatorname{ad}_{\overline{h}_{-1}}$ is a derivation acting trivially on the generators $\psi_{i},\phi_{i}$. The second equality follows from Lemma 4.8. The theorem now follows directly from Lemma 4.7. ∎
Remark 4.10.
The choice of the imaginary root vectors $\Theta_{s}$ in new Drinfeld presentation of the $q$-Onsager algebra in [LW21, §2.7] was suggested by the construction of the $\imath$Hall algebra of the projective line from [LRW23]. Earlier, a different set of root vectors $\acute{\Theta}_{s}$ (see (3.2)) was found in [BK20]. The corresponding Drinfeld presentation can be found in [LW21, §5.3]. Theorem 4.9 shows that the generators $\acute{\Theta}_{s}$, further renormalized by $(q-q^{-1})$, are indeed more natural from the point of view of representation theory, as already observed in Remark 3.2. Define
(4.85) $\grave{\Theta}_{s}=(q-q^{-1})\acute{\Theta}_{s}.$
Then Theorem 4.9 takes the following more elegant form.
Corollary 4.11.
Assuming $s_{0}=s_{1}=0$, we have
(4.86) $\boldsymbol{\grave{\Theta}}(z)\equiv\boldsymbol{\phi}(z^{-1})\boldsymbol{\psi}(Cz){\mod U_{+}}[\negthinspace[z]\negthinspace].$
5. Drinfeld rational fractions
5.1. Spectra of restricted representations
Let $\mathcal{H}$ (resp. $\mathcal{H}_{\mathcal{O}_{q}}$) be the commutative subalgebra of $U_{q}L\mathfrak{sl}_{2}$ (resp. $\mathcal{O}_{q}$) generated by elements $\phi_{i},\psi_{i}$ (resp. $\Theta_{i}$). Recall that every irreducible finite dimensional $U_{q}L\mathfrak{sl}_{2}$-module $V$ can be endowed with a ${\mathbb{Z}}_{\leq 0}$-grading (compatible with the ${\mathbb{Z}}$-grading on $U_{q}L\mathfrak{sl}_{2}$) such that the highest weight vector is in degree zero. Indeed, this is clear, in the case of evaluation representations, from their explicit description, and follows in the general case from the fact that every finite dimensional irreducible module is obtained as a tensor product of evaluation representations [CP91, Theorem 4.11]. This grading also induces a (vector space) filtration with $F_{i}(V)=\{v\in V\mid\deg v\geq i\}$. Since this filtration is preserved by the action of $\mathcal{H}$, we can consider the associated graded $\operatorname{gr}(V)$ as a $\mathcal{H}$-module.
The $\mathcal{O}_{q}$-module $\eta_{\mathbf{c},0}^{*}(V)$ has a decomposition into generalized $\mathcal{H}_{\mathcal{O}_{q}}$-eigenspaces. Let $w$ be generalized $\mathcal{H}_{\mathcal{O}_{q}}$-eigenvector and $\overline{w}$ its image in $\operatorname{gr}(V)$. Corollary 4.11 implies that $\overline{w}$ is an $\mathcal{H}$-eigenvector, and that the $\mathcal{H}_{\mathcal{O}_{q}}$-eigenvalues of $w$ can be computed using the $\mathcal{H}$-eigenvalues of $\overline{w}$. More precisely, in terms of generating series, the eigenvalue of $\boldsymbol{\grave{\Theta}}(z)$ on $w$ is the same as the $\boldsymbol{\phi}(z^{-1})\boldsymbol{\psi}(Cz)$-eigenvalue on $\overline{w}$. This fact has many interesting consequences.
5.2. Frenkel–Reshetikhin-type theorem
Let $w\in\eta_{\mathbf{c},0}^{*}(V)$ be a generalized $\mathcal{H}_{\mathcal{O}_{q}}$-eigenvector. Let $\boldsymbol{d}_{w}^{\pm}(z)\in{\mathbb{C}}[\negthinspace[z^{\pm 1}]\negthinspace]$ and $\boldsymbol{D}_{w}(z)\in{\mathbb{C}}[\negthinspace[z]\negthinspace]$ be defined by
(5.1) $\boldsymbol{\psi}(z)\cdot\overline{w}=\boldsymbol{d}_{w}^{+}(z)\overline{w},\quad\boldsymbol{\phi}(z)\cdot\overline{w}=\boldsymbol{d}_{w}^{-}(z)\overline{w},\quad\boldsymbol{\grave{\Theta}}(z)\cdot w=\boldsymbol{D}_{w}(z)w.$
By [FR99, Proposition 1], there exist polynomials $Q(z),R(z)\in{\mathbb{C}}[z]$ with constant term $1$ such that
(5.2) $\boldsymbol{d}_{w}^{\pm}(z)=q^{\deg Q-\deg R}\frac{Q(q^{-1}z)R(qz)}{Q(qz)R(q^{-1}z)}.$
Given a polynomial $P(z)\in{\mathbb{C}}[z]$ with constant term $1$, let $P^{\dagger}(z)$ be the polynomial with constant term $1$ whose roots are obtained from those of $P(z)$ via the transformation $a\mapsto C^{-1}a^{-1}$; and let $P^{*}(z)$ be the polynomial with constant term $1$ whose roots are the inverses of the roots of $P(z)$. Observe that
(5.3) $P(z)=\gamma_{P}z^{\deg P}P^{\dagger}(C^{-1}z^{-1}),\qquad P(z)=\gamma_{P}z^{\deg P}P^{*}(z^{-1}),$
where $\gamma_{P}$ is the product of the roots of $P^{*}(z)$.
Corollary 5.1.
The eigenvalues of $\boldsymbol{\grave{\Theta}}(z)$ on $w$ are of the form:
(5.4) $\boldsymbol{D}_{w}(z)=\frac{\mathbf{Q}(q^{-1}z)}{\mathbf{Q}(qz)}\frac{\mathbf{Q}^{\dagger}(qz)}{\mathbf{Q}^{\dagger}(q^{-1}z)},$
as elements of ${\mathbb{C}}[\negthinspace[z]\negthinspace]$, where $\mathbf{Q}(z)$ is a polynomial with constant term $1$. Explicitly,
$\mathbf{Q}(z)={Q(Cz)R^{*}(z)},\quad\mathbf{Q}^{\dagger}(z)={R(Cz)Q^{*}(z)}.$
Proof.
Corollary 4.11, together with the argument from §5.1, implies that
$\boldsymbol{D}_{w}(z)=\boldsymbol{d}_{w}^{+}(Cz)\boldsymbol{d}_{w}^{-}(z^{-1}).$
Using Frenkel and Reshetikhin’s formula (5.2), as well as (5.3), we get
(5.5) $\boldsymbol{D}_{w}(z)=$ $\ \frac{Q^{*}(qz)R^{*}(q^{-1}z)Q(Cq^{-1}z)R(Cqz)}{Q(Cqz)R(Cq^{-1}z)Q^{*}(q^{-1}z)R^{*}(qz)}$
(5.6) $=$ $\ \left(\frac{Q(Cq^{-1}z)R^{*}(q^{-1}z)}{Q(Cqz)R^{*}(qz)}\right)\left(\frac{R(Cqz)Q^{*}(qz)}{R(Cq^{-1}z)Q^{*}(q^{-1}z)}\right),$
as claimed. ∎
Remark 5.2.
It is interesting to compare Corollary 5.1 with the corresponding results for quantum affine algebras.
Firstly, we observe that formula (5.4) has an extra symmetry that is absent in (5.2). Namely, the polynomials $Q(z)$ and $R(z)$ are not related in any special way, unlike $\mathbf{Q}(z)$ and $\mathbf{Q}^{\dagger}(z)$.
Secondly, for quantum affine algebras, the case of the highest weight vector $v_{0}$ is special because $R(z)=1$ in (5.2). This special case was proven by Chari and Pressley in [CP91, Theorem 3.4], and the polynomial $Q(z)$ is called the Drinfeld polynomial. In contrast, in our case, both $\mathbf{Q}(z)$ and $\mathbf{Q}^{\dagger}(z)$ still appear in (5.4) (with opposite $q^{2}$-shifts) for $v_{0}$.
This fact admits the following interpretation. The notion of a highest weight, familiar from the representation theory of quantum affine algebras, is not directly applicable to modules restricted to the $q$-Onsager algebra. The main obstacle is the lack of canonical annihilation operators analogous to $x_{k}^{+}$. In fact, (4.23) implies that, in general, none of the Drinfeld generators $A_{k}$ of $\mathcal{O}_{q}$ annihilate the highest weight vector $v_{0}$.
Definition 5.3.
We will refer to the polynomial $\mathbf{Q}(z)$ from Corollary 5.1 as the Drinfeld polynomial associated to an $\mathcal{H}_{\mathcal{O}_{q}}$-eigenspace in $\eta_{\mathbf{c},0}^{*}(V)$.
5.3. Rational fractions
We will now present another interpretation of the eigenvalues series $\boldsymbol{D}_{w}(z)$, namely, as a ratio of rational functions with an additional symmetry.
Definition 5.4.
The Drinfeld rational fraction [^3] Note that the Drinfeld rational fraction is defined up to sign. This ambiguity can be resolved by choosing a branch of the square root function or considering ${\mathbb{Z}}/2{\mathbb{Z}}$-orbits of such functions. (DRF) associated to an $\mathcal{H}_{\mathcal{O}_{q}}$-eigenspace in $\eta^{*}_{\mathbf{c},0}(V)$ is
(5.7) $F(z)=$ $\ \gamma_{\mathbf{Q}}^{-1}C^{\frac{1}{2}\deg\mathbf{Q}}\frac{\mathbf{Q}(z)}{\mathbf{Q}^{\dagger}(z)}\in\mathbb{C}(z),$
Corollary 5.5.
The eigenvalues of $\boldsymbol{\grave{\Theta}}(z)$ on $w$ can be expressed as the following ratio of Drinfeld rational fractions:
(5.8) $\boldsymbol{D}_{w}(z)=\frac{F(q^{-1}z)}{F(qz)}.$
Moreover, the rational function $F(z)$:
  1. (1)
    does not have a pole or zero at $z=0$ or $z=\infty$,
  2. (2)
    satisfies the $C$-twisted unitarity property:
    (5.9) $F(qC^{-1}z^{-1})=F(q^{-1}z)^{-1}.$
Proof.
The identity (5.8) follows directly from the definition (5.7) and (5.5). Since, by [FR99, Proposition 1], neither $P_{V}(z)$ nor $P^{*}_{V}(z)$ has a zero at $z=0$, part (1) follows. Next, using (5.3) , we compute
(5.10) $F(qC^{-1}z^{-1})=$ $\ \gamma_{\mathbf{Q}}^{-1}C^{\frac{1}{2}\deg\mathbf{Q}}\frac{\mathbf{Q}(qC^{-1}z^{-1})}{\mathbf{Q}^{\dagger}(qC^{-1}z^{-1})}$
(5.11) $=$ $\ \gamma_{\mathbf{Q}}^{-1}C^{\frac{1}{2}\deg\mathbf{Q}}\frac{\gamma_{\mathbf{Q}}(qC^{-1}z^{-1})^{\deg\mathbf{Q}}\mathbf{Q}^{\dagger}(q^{-1}z)}{\mathbf{Q}^{\dagger}(qC^{-1}z^{-1})}$
(5.12) $=$ $\ (qC^{-\frac{1}{2}}z^{-1})^{\deg\mathbf{Q}}\frac{\mathbf{Q}(q^{-1}z)}{\mathbf{Q}^{\dagger}(qC^{-1}z^{-1})},$
(5.13) $F(q^{-1}z)^{-1}=$ $\ \gamma_{\mathbf{Q}}C^{-\frac{1}{2}\deg\mathbf{Q}}\frac{\mathbf{Q}^{\dagger}(q^{-1}z)}{\mathbf{Q}(q^{-1}z)}$
(5.14) $=$ $\ \gamma_{\mathbf{Q}}C^{-\frac{1}{2}\deg\mathbf{Q}}\frac{\mathbf{Q}^{\dagger}(q^{-1}z)}{\gamma_{\mathbf{Q}}(q^{-1}z)^{\deg\mathbf{Q}}\mathbf{Q}^{\dagger}(qC^{-1}z^{-1})}$
(5.15) $=$ $\ (qC^{-\frac{1}{2}}z^{-1})^{\deg\mathbf{Q}}\frac{\mathbf{Q}(q^{-1}z)}{\mathbf{Q}^{\dagger}(qC^{-1}z^{-1})},$
which yields (2). ∎
Remark 5.6.
Equation (5.8) has the form of a ratio involving a single $q^{-2}$-shift, familiar from the theory of Drinfeld polynomials. However, we emphasize that this kind of formula only holds in our case when polynomials are replaced by rational functions.
The twisted unitarity property can be construed as an analogue of the condition $u^{\deg P}P(u^{-1})=q^{-\deg P}P(uq^{2})$ in [GM10, Theorem 4.6], which appeared in the setting of the symplectic twisted $q$-Yangian.
Note that, by [Kol14, Theorem 11.7], the $q$-Onsager algebra corresponds to the orthogonal twisted $q$-Yangian of rank one. Therefore, the results of [GM10] are not applicable in our case. There is also a major difference in methodology as op. cit. uses the FRT presentation.
Remark 5.7.
The notion of a Drinfeld polynomial, in the context of the $q$-Onsager algebra, has already been introduced by Ito–Terwilliger in [IT08, IT10]. However, the definition in op. cit. is completely different from our definitions of Drinfeld polynomials and rational fractions. In particular, a priori, it has nothing to do with Lu–Wang’s new Drinfeld presentation, introduced only twelve years later.
On the other hand, the two concepts share many similarities. They both appear to play an important role in the classification of finite dimensional modules and behave nicely with respect to tensor products. It would be interesting to understand the precise connection between the two notions.
Remark 5.8.
The idea of replacing the RHS of $\boldsymbol{d}_{V}^{+}(z)=\frac{P_{V}(q^{-2}z)}{P_{V}(z)}$ with more general rational functions has already appeared in [HJ12] in the study of category $\mathcal{O}$ for $U_{q}\widehat{\mathfrak{b}}$, and in [Her23] in the study of shifted quantum affine algebras.
6. One-dimensional representations
The action of $\mathcal{O}_{q}^{\mathbf{c}}$ on a one-dimensional representation $V$ is determined by the two scalars by which the generators $A_{-1}$ and $A_{0}$ act on $V$. Therefore, given any pair of such scalars $t_{-1},t_{0}$, we can construct $V$ by restricting the trivial $U_{q}L\mathfrak{sl}_{2}$-representation via (2.20), with $s_{0}=q^{2}c_{0}t_{-1}$ and $s_{1}=t_{0}$. Let $\mathbb{C}_{\mathbf{c},\mathbf{s}}=\eta_{\mathbf{c},\mathbf{s}}^{*}(\mathbb{C})$ denote this restriction, and let $\xi_{\mathbf{c},\mathbf{s}}\colon\mathcal{O}_{q}\to\mathbb{C}$ be the character corresponding to the action of $\mathcal{O}_{q}$ on $\mathbb{C}_{\mathbf{c},\mathbf{s}}$, as in (2.23). Abbreviate
$\boldsymbol{D}_{\mathbf{c},\mathbf{s}}(z)=\boldsymbol{D}_{\mathbb{C}_{\mathbf{c},\mathbf{s}}}(z)=\xi_{\mathbf{c},\mathbf{s}}(\boldsymbol{\grave{\Theta}}(z)).$
Lemma 6.1.
We have
(6.1) $\boldsymbol{D}_{\mathbf{c},\mathbf{s}}(z)=\frac{q^{-1}(q-q^{-1})^{2}c_{1}^{-1}C(\alpha z+\beta(1+Cz^{2}))z+(1-Cz^{2})^{2}}{(1-Cz^{2})^{2}},$
where
$\alpha=C(q^{-2}c_{0}^{-1}s_{0})^{2}+s_{1}^{2},\quad\beta=q^{-2}c_{0}^{-1}s_{0}s_{1}.$
Proof.
Applying $\xi_{\mathbf{c},\mathbf{s}}$ to (3.4) and multiplying by $q-q^{-1}$, we get
(6.2) $\boldsymbol{D}_{\mathbf{c},\mathbf{s}}(z)=1+(1-Cz^{2})^{-1}c_{1}^{-1}Czq^{-1}(q-q^{-1})^{2}(\xi_{\mathbf{c},\mathbf{s}}(A_{-1})+\xi_{\mathbf{c},\mathbf{s}}(A_{0})z)\xi_{\mathbf{c},\mathbf{s}}(\boldsymbol{A}_{+}(z)).$
Relation (2.17) implies that
$\xi_{\mathbf{c},\mathbf{s}}(A_{2r})=C^{r}s_{1},\quad\xi_{\mathbf{c},\mathbf{s}}(A_{2r-1})=C^{r}q^{-2}c_{0}^{-1}s_{0},$
so
(6.3) $\xi_{\mathbf{c},\mathbf{s}}(\boldsymbol{A}_{+}(z))=(1-Cz^{2})^{-1}(s_{1}+Cq^{-2}c_{0}^{-1}s_{0}z).$
Substituting (6.3) into (6.2) yields the required formula. ∎
Proposition 6.2.
There exists a rational function $F_{\mathbf{c},\mathbf{s}}(z)\in\mathbb{C}(z)$ such that
$\boldsymbol{D}_{\mathbf{c},\mathbf{s}}(z)=\left(F_{\mathbf{c},\mathbf{s}}(z)F_{\mathbf{c},\mathbf{s}}(C^{-1}z^{-1})\right)^{-1}.$
This function is unique up to sign and transforming its poles under the map $a\mapsto C^{-1}a^{-1}$. Moreover:
  1. (1)
    $F_{\mathbf{c},\mathbf{s}}(z)=\pm 1$ if and only if $s_{0}=s_{1}=0$,
  2. (2)
    the numerator and denominator of $F_{\mathbf{c},\mathbf{s}}(z)$ have degree one if and only if $s_{0}\neq 0\neq s_{1}$ and $\frac{s_{0}}{s_{1}}=\pm\sqrt{\frac{c_{0}}{c_{1}}}$.
Proof.
Let
$G(z)=q^{-1}(q-q^{-1})^{2}c_{1}^{-1}C(\alpha z+\beta(1+Cz^{2}))z+(1-Cz^{2})^{2}=C^{2}\prod_{i=1}^{4}(z-\gamma_{i}).$
Then, by Lemma 6.1,
$\boldsymbol{D}_{\mathbf{c},\mathbf{s}}(z)=\frac{-C^{-1}G(z)z^{-2}}{(1-Cz^{2})(1-C^{-1}z^{-2})}.$
By the $C$-symmetry of $\boldsymbol{D}_{c,s}(z)$, we have
$G(z)z^{-2}=C^{2}G(C^{-1}z^{-1})z^{2}.$
Hence $\gamma_{3}=C^{-1}\gamma_{1}^{-1}$, $\gamma_{4}=C^{-1}\gamma_{2}^{-1}$ and
$G(z)=(\gamma_{1}\gamma_{2})^{-1}C^{2}(z-\gamma_{1})(z-\gamma_{2})(C^{-1}z^{-1}-\gamma_{1})(C^{-1}z^{-1}-\gamma_{1})z^{2}.$
Therefore
(6.4) $F_{\mathbf{c},\mathbf{s}}(z)=\pm i\frac{(\gamma_{k}\gamma_{l}C)^{1/2}(z^{2}-C^{-1})}{(z-\gamma_{k})(z-\gamma_{l})},$
where $k\in\{1,3\}$ and $l\in\{2,4\}$. This proves the existence and uniqueness of $F_{\mathbf{c},\mathbf{s}}(z)$, completing the proof of the main statement.
We now deduce the three additional statements. We will use the following equalities, obtained by considering the coefficients of the polynomial $G(z)$ on $z^{3}$ and $z^{2}$, respectively:
(6.5) $\tilde{\beta}=q^{-1}(q-q^{-1})^{2}c_{1}^{-1}\beta=$ $\ \gamma_{1}+\gamma_{2}+C^{-1}(\gamma_{1}^{-1}+\gamma_{2}^{-1}),$
(6.6) $\tilde{\alpha}=q^{-1}(q-q^{-1})^{2}c_{1}^{-1}C\alpha=$ $\ (\gamma_{1}C+\gamma_{1}^{-1})(\gamma_{2}C+\gamma_{2}^{-1})+4C.$
It follows directly from the definition (6.4) that $F_{\mathbf{c},\mathbf{s}}(z)=\pm 1$ if and only if $\gamma_{1}=C^{-1/2}$ and $\gamma_{2}=-C^{-1/2}$. By (6.5)–(6.6), this is the case if and only if $\alpha=\beta=0$, which is equivalent to $s_{0}=s_{1}=0$. This proves (1).
Similarly, it follows from (6.4) that $F_{\mathbf{c},\mathbf{s}}(z)$ is a ratio of polynomials of degree one if and only if $\gamma_{1}=\pm C^{-1/2}$ and $\gamma_{2}\neq\mp C^{-1/2}$. Without loss of generality, take $\gamma_{1}=C^{-1/2}$. Then we obtain the quadratic equations:
$\gamma_{2}^{2}+(2C^{-1/2}-\tilde{\beta})+C^{-1}=0,\quad\gamma_{2}^{2}+(2C^{-1/2}-\textstyle\frac{1}{2}C^{-3/2}\tilde{\alpha})+C^{-1}=0,$
which imply that $\alpha=2C^{1/2}\beta$, i.e.,
$C(q^{-2}c_{0}^{-1}s_{0})^{2}+s_{1}^{2}=2C^{1/2}q^{-2}c_{0}^{-1}s_{0}s_{1}.$
After rearranging, we get
$(c_{1}^{1/2}s_{0}-c_{0}^{1/2}s_{1})^{2}=0.$
This proves (2). ∎
Remark 6.3.
Note that, unlike the Drinfeld rational functions from §5.3, the functions $F_{\mathbf{c},\mathbf{s}}(z)$ are, in general, not twisted unitary (in the sense of (5.9)).
7. Multiplicative property
The goal of this section is to prove a more general version of the factorization theorem, where we drop the assumption that $s_{0}=s_{1}=0$. To do this, we prove a multiplicative property for the series $\boldsymbol{\grave{\Theta}}(z)$ with respect to actions on one-dimensional representations. The key ingredients are formulas for the coproduct of $\boldsymbol{A}_{+}(z)$ and $\boldsymbol{\grave{\Theta}}(z)$, modulo certain ideals, established in §7.17.2 below. Once again, we emphasize that the parameters $s_{0},s_{1}$ can be arbitrary throughout §7.
7.1. Coproduct formula for $\boldsymbol{A}_{+}(z)$
It follows directly from (2.1) and (2.21) that
$\Delta_{\mathbf{c},\mathbf{s}}(A_{-1})=1\otimes\eta_{\mathbf{c},0}(A_{-1})+\eta_{\mathbf{c},\mathbf{s}}(A_{-1})\otimes K,\quad\Delta_{\mathbf{c},\mathbf{s}}(A_{0})=1\otimes\eta_{\mathbf{c},0}(A_{0})+\eta_{\mathbf{c},\mathbf{s}}(A_{0})\otimes K^{-1}.$
For the sake of readability, we abbreviate $Y\otimes 1=\eta_{\mathbf{c},\mathbf{s}}(Y)\otimes 1$ and $1\otimes Y=1\otimes\eta_{\mathbf{c},0}(Y)$, for $Y\in\mathcal{O}_{q}^{\mathbf{c}}$. Relation (3.9) implies that
$\Delta(\overline{H}_{1})=1\otimes\overline{H}_{1}+\overline{H}_{1}\otimes 1-(q-q^{-1})A_{0}\otimes x^{+}_{-1}-C(q-q^{-1})A_{-1}\otimes x^{+}_{0}.$
We will denote the four summands on the RHS above as $Q_{1},\cdots,Q_{4}$, respectively. We also abbreviate
(7.1) $S^{r}_{1}=$ $\ 1\otimes A_{r},$
(7.2) $S^{r}_{2}=$ $\ A_{0}\otimes\phi_{-r},$
(7.3) $S^{r}_{3}=$ $\ \sum_{i=0}^{r-1}A_{r-i}\otimes\phi_{-i}.$
We write sums of the terms above as, e.g., $Q_{1+2},S^{r}_{2+3},S^{r}_{1+2+3}$, etc. Throughout this subsection the symbol $\equiv$” denotes equalities holding modulo $1\otimes U_{\geq 2}$.
Lemma 7.1.
We have
$S^{r+1}_{1+2+3}\equiv[Q_{1+2},S^{r}_{1+2+3}]+CS^{r-1}_{1+2+3}+[Q_{3+4},S^{r}_{1}]\mod 1\otimes U_{\geq 2}.$
Proof.
We use relation (4.21), i.e.,
(7.4) $A_{r+1}=[\overline{H}_{1},A_{r}]+CA_{r-1},$
throughout.
Now assume $r\geq 2$. Using (7.4) and Lemma 4.5, we get
(7.5) $[Q_{1},S^{r}_{1}]+CS^{r-1}_{1}=$ $\ S^{r+1}_{1},$
(7.6) $[Q_{3},S^{r}_{1}]\equiv$ $\ A_{0}\otimes(q-q^{-1})[x^{+}_{-1},x^{-}_{-r}]$
(7.7) $=$ $\ S^{r+1}_{2}.$
Next, using (7.4) repeatedly, we see that
$[Q_{2},S^{r}_{3}]+CS^{r-1}_{2+3}=\sum_{i=0}^{r-1}A_{r+1-i}\otimes\phi_{-i}.$
Moreover,
(7.8) $[Q_{4},S^{r}_{1}]+[Q_{2},S^{r}_{2}]=$ $\ -CA_{-1}\otimes(q-q^{-1})[x^{+}_{0},x^{-}_{-r}]+[\overline{H}_{1},A_{0}]\otimes\phi_{-r}$
(7.9) $=$ $\ ([\overline{H}_{1},A_{0}]+CA_{-1})\otimes\phi_{-r}=A_{1}\otimes\phi_{-r}.$
Hence
$[Q_{2},S^{r}_{2+3}]+CS^{r-1}_{2+3}+[Q_{4},S^{r}_{1}]=S^{r+1}_{3}.$
Next, observe that
$[Q_{1},S^{r}_{2+3}]\equiv[Q_{2},S^{r}_{1}]\equiv 0.$
Summing up our calculations so far, we, therefore, obtain
$[Q_{1+2},S^{r}_{1+2+3}]+CS^{r-1}_{1+2+3}+[Q_{3+4},S^{r}_{1}]\equiv S^{r+1}_{1+2+3},$
as claimed. ∎
Write
$\overline{\Delta}_{\mathbf{c},\mathbf{s}}=(\widetilde{\xi}_{\mathbf{c},\mathbf{s}}\otimes\operatorname{id})\circ\Delta_{\mathbf{c},0},$
where $\widetilde{\xi}_{\mathbf{c},\mathbf{s}}$ is as in (2.23). From now on all the elements on the LHS of the tensor symbol are understood as images under $\xi_{\mathbf{c},\mathbf{s}}$. For example, $A_{0}\otimes\phi_{-r}=\xi_{\mathbf{c},\mathbf{s}}(A_{0})\otimes\phi_{-r}$, etc. In particular, any two elements on the LHS of the tensor symbol commute. Abbreviate
$\kappa=-(q-q^{-1})\big{(}\xi_{\mathbf{c},\mathbf{s}}(A_{0})+C\xi_{\mathbf{c},\mathbf{s}}(A_{-1})\operatorname{ad}_{\overline{h}_{1}}\big{)}$
so that
$Q_{3+4}=1\otimes\kappa x^{+}_{-1}.$
Also abbreviate
(7.10) $S^{r}_{4}=\sum_{\begin{subarray}{c}i+j+k=r;\\ i\geq 1;j,k\geq 0\end{subarray}}A_{k}\otimes\kappa\mathbbm{h}_{i-1}[x^{+}_{-1},\phi_{-j}],$
where $\mathbbm{h}_{i}$ denotes the $i$-th degree complete homogeneous symmetric polynomial in $-\operatorname{ad}_{\overline{h}_{-1}}$ and $C\operatorname{ad}_{\overline{h}_{1}}$, so that $\sum_{i=0}^{\infty}\mathbbm{h}_{i}z^{i}=\Omega^{-1},$ with $\Omega$ as in (4.29).
Lemma 7.2.
We have
$S^{r+1}_{4}\equiv[Q_{1+2+3+4},S^{r}_{4}]+CS^{r-1}_{4}+[Q_{3+4},S^{r}_{2+3}]\mod 1\otimes U_{\geq 2}.$
Proof.
Clearly
$[Q_{2+3+4},S_{4}^{r}]\equiv 0.$
Next, observe that
(7.11) $[\overline{H}_{1},\mathbbm{h}_{i-1}]+C\mathbbm{h}_{i-2}\equiv$ $\ (C\operatorname{ad}_{\overline{h}_{1}}-\operatorname{ad}_{\overline{h}_{-1}})\mathbbm{h}_{i-1}+C\mathbbm{h}_{i-2}-C\operatorname{ad}_{\overline{h}_{1}}\operatorname{ad}_{\overline{h}_{-1}}\mathbbm{h}_{i-2}=\mathbbm{h}_{i}.$
Hence
(7.12) $[Q_{1},S^{r}_{4}]+CS^{r-1}_{4}+[Q_{3+4},S^{r}_{2+3}]\equiv S^{r+1}_{4},$
as claimed. ∎
Proposition 7.3.
For each $r\geq 0$,
$\overline{\Delta}_{\mathbf{c},\mathbf{s}}(A_{r})\equiv S^{r}_{1+2+3+4}\mod 1\otimes U_{\geq 2}.$
Hence
(7.13) $\overline{\Delta}_{\mathbf{c},\mathbf{s}}(\boldsymbol{A}_{+}(z))\equiv$ $\ 1\otimes\boldsymbol{A}_{+}(z)+\boldsymbol{A}_{+}(z)\otimes\boldsymbol{\phi}(z^{-1})+\boldsymbol{A}_{+}(z)\otimes\kappa\Gamma,$
where
$\Gamma=(1-q^{2}\operatorname{ad}_{\overline{h}_{-1}}z)^{-1}(1-C\operatorname{ad}_{\overline{h}_{1}}z)^{-1}(q^{2}-1)\boldsymbol{\phi}(z^{-1})x^{+}_{-1}z.$
Proof.
The first statement follows by induction, using (7.4) and Lemmas 7.17.2. It also follows directly from the definition of $S^{r}_{k}$ that
$\sum_{r\geq 0}S^{r}_{1}=1\otimes\boldsymbol{A}_{+}(z),\qquad\sum_{r\geq 0}S^{r}_{2+3}=\boldsymbol{A}_{+}(z)\otimes\boldsymbol{\phi}(z^{-1}).$
Moreover,
(7.14) $\sum_{r\geq 1}S^{r}_{4}z^{r}=$ $\ \boldsymbol{A}_{+}(z)\otimes z\kappa\Omega^{-1}[x^{+}_{-1},\boldsymbol{\phi}(z^{-1})]=\boldsymbol{A}_{+}(z)\otimes\kappa\Gamma,$
where the last equality follows from Lemma 4.2. ∎
We will also use the notation
$\widetilde{\Gamma}=(1-q^{2}\operatorname{ad}_{\overline{h}_{-1}}z)^{-1}(1-C\operatorname{ad}_{\overline{h}_{1}}z)^{-1}(q^{2}-1)x^{+}_{-1}z$
so that $\Gamma=\boldsymbol{\phi}(z^{-1})\widetilde{\Gamma}$.
7.2. Coproduct formula for $\boldsymbol{\grave{\Theta}}(z)$
Throughout this subsection the symbol $\equiv$” denotes equalities holding modulo $1\otimes U_{+}$.
Proposition 7.4.
We have
$\overline{\Delta}_{\mathbf{c},\mathbf{s}}(\boldsymbol{\grave{\Theta}}(z))\equiv\xi_{\mathbf{c},\mathbf{s}}(\boldsymbol{\grave{\Theta}}(z))\cdot\eta_{\mathbf{c},0}(\boldsymbol{\grave{\Theta}}(z))\mod{1\otimes U_{+}}[\negthinspace[z]\negthinspace].$
Proof.
We use (3.12) and (7.13) throughout. Let us denote the three summands on the RHS of (7.13) as $R_{1},R_{2},R_{3}$, respectively. Then
(7.15) $(1-Cz^{2})\overline{\Delta}_{\mathbf{c},\mathbf{s}}(\Theta_{0})\ +z^{2}c_{1}^{-1}C\big{(}z^{-1}[1\otimes A_{-1},R_{1}]_{q^{-2}}-q^{-2}[$ $1\otimes A_{0},R_{1}]_{q^{2}}\big{)}=$
(7.16) $=$ $\ 1\otimes(1-q^{-2}Cz^{2})\boldsymbol{\Theta}(z),$
and
(7.17) $z^{2}c_{1}^{-1}C\big{(}$ $z^{-1}[A_{-1}\otimes K,R_{1}]_{q^{-2}}-q^{-2}[A_{0}\otimes K^{-1},R_{1}]_{q^{2}}\big{)}\equiv$
(7.18) $\equiv$ $\ z^{2}c_{1}^{-1}C\left(z^{-1}A_{-1}\otimes[K,\boldsymbol{x}^{-}(z^{-1})]_{q^{-2}}-q^{-2}A_{0}\otimes[K^{-1},\boldsymbol{x}^{-}(z^{-1})]_{q^{2}}\right)$
(7.19) $\equiv$ $\ 0,$
because the terms on the RHS of the tensor symbols vanish. This implies that
(7.20) $(1-Cz^{2})\overline{\Delta}_{\mathbf{c},\mathbf{s}}(\Theta_{0})\ +z^{2}c_{1}^{-1}C\big{(}z^{-1}[\overline{\Delta}_{\mathbf{c},\mathbf{s}}(A_{-1}),R_{1}]_{q^{-2}}-$ $q^{-2}[\overline{\Delta}_{\mathbf{c},\mathbf{s}}(A_{0}),R_{1}]_{q^{2}}\big{)}\equiv$
(7.21) $\equiv$ $\ 1\otimes(1-q^{-2}Cz^{2})\boldsymbol{\Theta}(z).$
Next, it follows from Lemma 4.3 that
(7.22) $z^{2}c_{1}^{-1}C\big{(}z^{-1}[$ $1\otimes A_{-1},R_{2}]_{q^{-2}}-q^{-2}[1\otimes A_{0},R_{2}]_{q^{2}}\big{)}\equiv$
(7.23) $\equiv$ $\ z^{2}c_{1}^{-1}C\boldsymbol{A}_{+}(z)\otimes\left(z^{-1}[x^{-}_{1},\boldsymbol{\phi}(z^{-1})]_{q^{-2}}-q^{-2}[x^{-}_{0},\boldsymbol{\phi}(z^{-1})]_{q^{2}}\right)$
(7.24) $\equiv$ $\ 0,$
and it follows directly from the definitions that
(7.25) $z^{2}c_{1}^{-1}C\big{(}z^{-1}[$ $A_{-1}\otimes K,R_{2}]_{q^{-2}}-q^{-2}[A_{0}\otimes K^{-1},R_{2}]_{q^{2}}\big{)}\equiv$
(7.26) $\equiv$ $\ z^{2}c_{1}^{-1}C(1-q^{-2})\boldsymbol{A}_{+}(z)\otimes\left(z^{-1}\xi_{\mathbf{c},\mathbf{s}}(A_{-1})K+\xi_{\mathbf{c},\mathbf{s}}(A_{0})K^{-1}\right)\boldsymbol{\phi}(z^{-1}).$
This implies that
(7.27) $z^{2}c_{1}^{-1}C\big{(}z^{-1}[$ $\overline{\Delta}_{\mathbf{c},\mathbf{s}}(A_{-1}),R_{2}]_{q^{-2}}-q^{-2}[\overline{\Delta}_{\mathbf{c},\mathbf{s}}(A_{0}),R_{2}]_{q^{2}}\big{)}\equiv$
(7.28) $\equiv$ $\ z^{2}c_{1}^{-1}C(1-q^{-2})\boldsymbol{A}_{+}(z)\otimes\left(z^{-1}\xi_{\mathbf{c},\mathbf{s}}(A_{-1})K+\xi_{\mathbf{c},\mathbf{s}}(A_{0})K^{-1}\right)\boldsymbol{\phi}(z^{-1}).$
Moreover,
(7.29) $z^{2}c_{1}^{-1}C\big{(}z^{-1}[\overline{\Delta}_{\mathbf{c},\mathbf{s}}(A_{-1}),$ $\ R_{3}]_{q^{-2}}-q^{-2}[\overline{\Delta}_{\mathbf{c},\mathbf{s}}(A_{0}),R_{3}]_{q^{2}}\big{)}\equiv$
(7.30) $\equiv$ $\ z^{2}c_{1}^{-1}C\boldsymbol{A}_{+}(z)\otimes\big{(}z^{-1}[x^{-}_{1},\kappa\Gamma]_{q^{-2}}-q^{-2}[x^{-}_{0},\kappa\Gamma]_{q^{2}}\big{)}$
(7.31) $\equiv$ $\ -q^{-2}zc_{1}^{-1}C\boldsymbol{A}_{+}(z)\otimes\boldsymbol{\phi}(z^{-1})[\kappa\widetilde{\Gamma},(1+q^{2}\operatorname{ad}_{\overline{h}_{-1}}z)x^{-}_{1}]$
(7.32) $\equiv$ $\ -q^{-2}zc_{1}^{-1}C\boldsymbol{A}_{+}(z)\otimes\boldsymbol{\phi}(z^{-1})[(1-q^{2}\operatorname{ad}_{\overline{h}_{-1}}z)\kappa\widetilde{\Gamma},x^{-}_{1}]$
(7.33) $\equiv$ $\ q^{-1}z^{2}(q-q^{-1})^{2}c_{1}^{-1}C\boldsymbol{A}_{+}(z)\otimes\boldsymbol{\phi}(z^{-1})\cdot$
(7.34) $\cdot\big{[}(1-C\operatorname{ad}_{\overline{h}_{1}}z)^{-1}(\xi_{\mathbf{c},\mathbf{s}}(A_{0})x^{+}_{-1}+C\xi_{\mathbf{c},\mathbf{s}}(A_{-1})x^{+}_{0}),x^{-}_{1}\big{]},$
where Lemma 4.3 was used in the third equality. Let us calculate the term in the last line above. Observe that
(7.35) $(1-C\operatorname{ad}_{\overline{h}_{1}}z)^{-1}(\xi_{\mathbf{c},\mathbf{s}}(A_{0})x^{+}_{-1}\ +$ $\ C\xi_{\mathbf{c},\mathbf{s}}(A_{-1})x^{+}_{0})=$
(7.36) $=$ $\ C(\xi_{\mathbf{c},\mathbf{s}}(A_{-1})+\xi_{\mathbf{c},\mathbf{s}}(A_{0})z)\boldsymbol{x}^{+}_{\geq 0}(Cz)+\xi_{\mathbf{c},\mathbf{s}}(A_{0})x^{+}_{-1}.$
Hence
(7.37) $(q-q^{-1})\big{[}$ $(1-C\operatorname{ad}_{\overline{h}_{1}}z)^{-1}(\xi_{\mathbf{c},\mathbf{s}}(A_{0})x^{+}_{-1}+C\xi_{\mathbf{c},\mathbf{s}}(A_{-1})x^{+}_{0}),x^{-}_{1}\big{]}=$
(7.38) $=$ $\ z^{-1}(\xi_{\mathbf{c},\mathbf{s}}(A_{-1})+\xi_{\mathbf{c},\mathbf{s}}(A_{0})z)\boldsymbol{\psi}_{\geq 1}(Cz)+\xi_{\mathbf{c},\mathbf{s}}(A_{0})(K-K^{-1})$
(7.39) $=$ $\ (\xi_{\mathbf{c},\mathbf{s}}(A_{-1})z^{-1}+\xi_{\mathbf{c},\mathbf{s}}(A_{0}))\boldsymbol{\psi}(Cz)-\xi_{\mathbf{c},\mathbf{s}}(A_{-1})z^{-1}K-\xi_{\mathbf{c},\mathbf{s}}(A_{0})K^{-1}.$
It follows that
(7.40) $z^{2}c_{1}^{-1}C\big{(}z^{-1}[\overline{\Delta}_{\mathbf{c},\mathbf{s}}(A_{-1}),R_{3}]_{q^{-2}}\ -$ $\ q^{-2}[\overline{\Delta}_{\mathbf{c},\mathbf{s}}(A_{0}),R_{3}]_{q^{2}}\big{)}\equiv$
(7.41) $\equiv$ $\ q^{-1}z^{2}(q-q^{-1})c_{1}^{-1}C\boldsymbol{A}_{+}(z)\otimes\boldsymbol{\phi}(z^{-1})\cdot$
(7.42) $\{(\xi_{\mathbf{c},\mathbf{s}}(A_{-1})z^{-1}+\xi_{\mathbf{c},\mathbf{s}}(A_{0}))\boldsymbol{\psi}(Cz)$
(7.43) $-\xi_{\mathbf{c},\mathbf{s}}(A_{-1})z^{-1}K-\xi_{\mathbf{c},\mathbf{s}}(A_{0})K^{-1}\}.$
The summand corresponding to the last line above cancels out (7.28). Therefore, collecting (7.21)–(7.40) yields
(7.44) $(1-q^{-2}Cz^{2})\overline{\Delta}_{\mathbf{c},\mathbf{s}}(\boldsymbol{\Theta}(z))\equiv$ $\ 1\otimes(1-q^{-2}Cz^{2})\boldsymbol{\Theta}(z)$
(7.45) $+z(1-q^{-2})c_{1}^{-1}C(A_{-1}+A_{0}z)\boldsymbol{A}_{+}(z)\otimes\boldsymbol{\phi}(z^{-1})\boldsymbol{\psi}(Cz).$
Using Corollary 4.11, this simplifies to
(7.46) $(1-q^{-2}Cz^{2})\overline{\Delta}_{\mathbf{c},\mathbf{s}}(\boldsymbol{\Theta}(z))\equiv$ $\ 1\otimes(1-q^{-2}Cz^{2})\boldsymbol{\Theta}(z)$
(7.47) $+z(1-q^{-2})c_{1}^{-1}C(A_{-1}+A_{0}z)\boldsymbol{A}_{+}(z)\otimes\boldsymbol{\grave{\Theta}}(z).$
The proposition now follows directly from (6.2). ∎
7.3. General factorization theorem
We can now prove the main result of this section.
Theorem 7.5.
For any $\mathbf{s}\in{\mathbb{C}}^{2}$, the following factorization and coproduct formulae hold:
  1. (1)
    generalized factorization:
    $\eta_{\mathbf{c},\mathbf{s}}(\boldsymbol{\grave{\Theta}}(z))\equiv\xi_{\mathbf{c},\mathbf{s}}(\boldsymbol{\grave{\Theta}}(z))\cdot\eta_{\mathbf{c},0}(\boldsymbol{\grave{\Theta}}(z))\quad\mod{(1\otimes U_{+})}[\negthinspace[z]\negthinspace],$
  2. (2)
    coproduct formula - the series $\boldsymbol{\grave{\Theta}}(z)$ is “group-like”, i.e.,
    $\Delta_{\mathbf{c},\mathbf{s}}(\boldsymbol{\grave{\Theta}}(z))\equiv\eta_{\mathbf{c},\mathbf{s}}(\boldsymbol{\grave{\Theta}}(z))\otimes\eta_{\mathbf{c},0}(\boldsymbol{\grave{\Theta}}(z))\quad\mod{(U_{-}\otimes U_{+}+U_{+}\otimes U_{-})}[\negthinspace[z]\negthinspace].$
Proof.
The first part is a direct consequence of Proposition 7.4 and (2.24). For the second part, we have
(7.48) $\Delta_{\mathbf{c},\mathbf{s}}(\boldsymbol{\grave{\Theta}}(z))\equiv$ $\ \Delta\circ\eta_{\mathbf{c},\mathbf{s}}(\boldsymbol{\grave{\Theta}}(z))$
(7.49) $\equiv$ $\ \xi_{\mathbf{c},\mathbf{s}}(\boldsymbol{\grave{\Theta}}(z))\cdot\Delta\circ\eta_{\mathbf{c},0}(\boldsymbol{\grave{\Theta}}(z))$
(7.50) $\equiv$ $\ \xi_{\mathbf{c},\mathbf{s}}(\boldsymbol{\grave{\Theta}}(z))\cdot\Delta_{\mathbf{c},0}(\boldsymbol{\grave{\Theta}}(z))$
(7.51) $\equiv$ $\ \xi_{\mathbf{c},\mathbf{s}}(\boldsymbol{\grave{\Theta}}(z))\cdot\eta_{\mathbf{c},0}(\boldsymbol{\grave{\Theta}})\otimes\eta_{\mathbf{c},0}(\boldsymbol{\grave{\Theta}})$
(7.52) $\equiv$ $\ \eta_{\mathbf{c},\mathbf{s}}(\boldsymbol{\grave{\Theta}})\otimes\eta_{\mathbf{c},0}(\boldsymbol{\grave{\Theta}}),$
where the second and fifth equalities follow from part (1) of the present theorem, and the fourth equality from Theorem 4.9 and the group-like property of the series $\boldsymbol{\psi}(z),\boldsymbol{\phi}(z)$ (see, e.g., [FR99, Lemma 1], originally proven in [Dam98]). ∎
As a consequence, we also obtain a generalization of Corollary 5.5. Let $V$ be an irreducible finite dimensional $U_{q}L\mathfrak{sl}_{2}$-module and $\eta_{\mathbf{c},\mathbf{s}}^{*}(V)$ its restriction to $\mathcal{O}_{q}^{\mathbf{c}}$ via the embedding (2.20), with $s_{0},s_{1}$ arbitrary. Let $w\in\eta_{\mathbf{c},\mathbf{s}}^{*}(V)$ be a generalized $\mathcal{H}_{\mathcal{O}_{q}}$-eigenvector. Let $\boldsymbol{D}_{\eta_{\mathbf{c},\mathbf{s}}^{*}(V),w}(z)\in{\mathbb{C}}[\negthinspace[z]\negthinspace]$ be defined by $\boldsymbol{\grave{\Theta}}(z)\cdot w=\boldsymbol{D}_{\eta_{\mathbf{c},\mathbf{s}}^{*}(V),w}(z)w$.
The following result says that the series of the eigenvalues of the operators $\grave{\Theta}_{r}$ on $w$ factorizes into the the corresponding series for the standard restriction $\eta_{\mathbf{c},0}^{*}(V)$ and the one-dimensional representation ${\mathbb{C}}_{\mathbf{c},\mathbf{s}}$.
Corollary 7.6.
The series $\boldsymbol{D}_{\eta_{\mathbf{c},\mathbf{s}}^{*}(V)}(z)$ can be expressed as the product:
$\boldsymbol{D}_{\eta_{\mathbf{c},\mathbf{s}}^{*}(V),w}(z)=\boldsymbol{D}_{\mathbf{c},\mathbf{s}}(z)\cdot\boldsymbol{D}_{\eta_{\mathbf{c},0}^{*}(V),w}(z).$
Proof.
The statement follows directly from Theorem 7.5. ∎
8. Affine QSP of split type
8.1. Quantum affine algebras
Let $\mathbb{I}_{0}=\{1,\cdots,N\}$ and $\mathbb{I}=\mathbb{I}_{0}\cup\{0\}$. Let ${\mathfrak{g}}$ be a simple Lie algebra of type $\mathsf{ADE}$ with Cartan matrix $(a_{ij})_{i,j\in\mathbb{I}_{0}}$, and $\widehat{{\mathfrak{g}}}$ the corresponding untwisted affine Lie algebra with affine Cartan matrix $(a_{ij})_{i,j\in\mathbb{I}}$. We use standard notations and conventions regarding root systems, weight lattices, Weyl groups, etc., as in, e.g., [LW21, §3.1]. In particular, we let $\alpha_{i}$ ($i\in\mathbb{I}$) denote the simple roots of $\widehat{{\mathfrak{g}}}$; let $P$ and $Q$ denote the weight and root lattices of ${\mathfrak{g}}$, respectively; and let $\omega_{i}\in P$ ($i\in\mathbb{I}_{0}$) be the fundamental weights of ${\mathfrak{g}}$.
The quantum affine algebra $U_{q}\widehat{\mathfrak{g}}$ is the algebra with generators $e_{i}^{\pm},K_{i}^{\pm 1}$ $(i\in\mathbb{I})$ and relations:
(8.1) $K_{i}K_{i}^{-1}$ $=K_{i}^{-1}K_{i}=1,$
(8.2) $K_{i}K_{j}$ $=K_{j}K_{i},$
(8.3) $K_{i}e_{j}^{\pm}$ $=q^{\pm a_{ij}}e_{j}^{\pm}K_{i},$
(8.4) $[e_{i}^{+},e_{j}^{-}]$ $=\delta_{ij}\frac{K_{i}-K_{i}^{-1}}{q-q^{-1}},$
(8.5) $\mathsf{Serre}_{ij}(e_{i}^{\pm},e_{j}^{\pm})$ $=0\qquad(i\neq j),$
where
[…]
We also write
$E_{i}=e_{i}^{+},\quad F_{i}=e_{i}^{-}.$
We refer to the presentation above as the Kac–Moody presentation of $U_{q}\widehat{\mathfrak{g}}$. The quantum loop algebra $U_{q}L\mathfrak{g}$ is the quotient of $U_{q}\widehat{\mathfrak{g}}$ by the ideal generated by the central element $K_{\delta}-1$, where $K_{\delta}$ is defined as in, e.g., [Bec94, §1.3].
Let us recall the “new” Drinfeld presentation of the quantum loop algebra $U_{q}L\mathfrak{g}$. By [Dri87a, Bec94], $U_{q}L\mathfrak{g}$ is isomorphic to the algebra generated by $x^{\pm}_{i,k},h_{i,l},K^{\pm 1}_{i}$, where $k\in{\mathbb{Z}}$, $l\in{\mathbb{Z}}-\{0\}$ and $i\in\mathbb{I}_{0}$, subject to the following relations:
(8.6) $K_{i}K_{i}^{-1}=$ $\ K_{i}^{-1}K_{i}=1,$
(8.7) $K_{i}K_{j}=$ $\ K_{j}K_{i},$
(8.8) $[h_{i,k},h_{j,l}]=$ $\ 0,$
(8.9) $K_{i}h_{j,k}=$ $\ h_{j,k}K_{i},$
(8.10) $K_{i}x^{\pm}_{j,k}=$ $\ q^{\pm a_{ij}}x^{\pm}_{j,k}K_{i},$
(8.11) $[h_{i,k},x^{\pm}_{j,l}]=$$\pm\textstyle\frac{[k\cdot a_{ij}]}{k}x^{\pm}_{j,k+l},$
(8.12) $x^{\pm}_{i,k+1}x^{\pm}_{j,l}-q^{\pm a_{ij}}x^{\pm}_{j,l}x^{\pm}_{i,r+1}=$ $\ q^{\pm a_{ij}}x^{\pm}_{i,k}x^{\pm}_{j,l+1}-x^{\pm}_{j,l+1}x^{\pm}_{i,k},$
(8.13) $[x^{+}_{i,k},x^{-}_{j,l}]=$ $\ \delta_{ij}\textstyle\frac{1}{q-q^{-1}}(\psi_{i,k+l}-\phi_{i,k+l}),$
[…]
where $\operatorname{Sym}_{k_{1},\cdots,k_{r}}$ denotes symmetrization with respect to the indices $k_{1},\cdots,k_{r}$ and
$\boldsymbol{\phi}^{\pm}_{i}(z)=\sum_{k=0}^{\infty}\phi^{\pm}_{i,\pm k}z^{\pm k}=K_{i}^{\pm 1}\exp\left(\pm(q-q^{-1})\sum_{k=1}^{\infty}h_{i,\pm k}z^{\pm k}\right).$
We will also consider the following version of the quantum loop algebra
$U_{q}^{\mathbb{K}}L\mathfrak{g}=U_{q}L\mathfrak{g}\otimes_{{\mathbb{C}}}{\mathbb{C}}[\mathbb{K}_{j}^{\pm 1}\mid j\in\mathbb{I}]$
with extended scalars, matching the central elements in $\mathcal{O}_{q}(\widehat{{\mathfrak{g}}})$ (see (8.41)). We consider $U_{q}^{\mathbb{K}}L\mathfrak{g}$ as a $Q$-graded algebra with
$\operatorname{deg}^{\mathrm{Dr}}x^{\pm}_{i,k}=\pm\alpha_{i},\quad\operatorname{deg}^{\mathrm{Dr}}h_{i,k}=\operatorname{deg}^{\mathrm{Dr}}\mathbb{K}_{i}=0\qquad(i\in\mathbb{I}_{0}).$
We will write $\operatorname{deg}^{\mathrm{Dr}}_{i}$ for degree along the simple root $\alpha_{i}$. Let $Q_{i}$ be the sublattice of $Q$ spanned by all the simple roots except $\alpha_{i}$. Let $U_{d_{i}=r,+}$ be the subalgebra of $U_{q}^{\mathbb{K}}L\mathfrak{g}$ spanned by elements which are both: (i) of degree $r$ along $\alpha_{i}$, and (ii) positive $Q_{i}$-degree, i.e., non-negative along all the simple roots and positive along at least one simple root in $Q_{i}$. That is,
(8.14) $U_{d_{i}=r,+}=$ $\ \langle y\in U_{q}^{\mathbb{K}}L\mathfrak{g}\mid\operatorname{deg}^{\mathrm{Dr}}_{i}y=r,\ \forall j\neq i\in\mathbb{I}_{0}\colon\operatorname{deg}^{\mathrm{Dr}}_{j}y\geq 0,\ \exists j\neq i\colon\operatorname{deg}^{\mathrm{Dr}}_{j}y>0\rangle,$
(8.15) $U_{d_{i}\geq r,+}=$ $\ \langle y\in U_{q}^{\mathbb{K}}L\mathfrak{g}\mid\operatorname{deg}^{\mathrm{Dr}}_{i}y\geq r,\ \forall j\neq i\in\mathbb{I}_{0}\colon\operatorname{deg}^{\mathrm{Dr}}_{j}y\geq 0,\ \exists j\neq i\colon\operatorname{deg}^{\mathrm{Dr}}_{j}y>0\rangle,$
We will also need the following subalgebras of $U_{q}^{\mathbb{K}}L\mathfrak{g}$:
(8.17) $U_{+}=$ $\ \langle y\in U_{q}^{\mathbb{K}}L\mathfrak{g}\mid\forall j\in\mathbb{I}_{0}\colon\operatorname{deg}^{\mathrm{Dr}}_{j}y\geq 0,\ \exists j\colon\operatorname{deg}^{\mathrm{Dr}}_{j}y>0\rangle,$
(8.18) $U_{-}=$ $\ \langle y\in U_{q}^{\mathbb{K}}L\mathfrak{g}\mid\forall j\in\mathbb{I}_{0}\colon\operatorname{deg}^{\mathrm{Dr}}_{j}y\leq 0,\ \exists j\colon\operatorname{deg}^{\mathrm{Dr}}_{j}y<0\rangle,$
(8.19) $U_{\neq i,+}=$ $\ \langle y\in U_{q}^{\mathbb{K}}L\mathfrak{g}\mid\forall j\neq i\in\mathbb{I}_{0}\colon\operatorname{deg}^{\mathrm{Dr}}_{j}y\geq 0,\ \exists j\neq i\colon\operatorname{deg}^{\mathrm{Dr}}_{j}y>0\rangle,$
(8.20) $U_{\neq i,-}=$ $\ \langle y\in U_{q}^{\mathbb{K}}L\mathfrak{g}\mid\forall j\neq i\in\mathbb{I}_{0}\colon\operatorname{deg}^{\mathrm{Dr}}_{j}y\leq 0,\ \exists j\neq i\colon\operatorname{deg}^{\mathrm{Dr}}_{j}y<0\rangle,$
(8.21) $U_{i+,0}=$ $\ \langle y\in U_{q}^{\mathbb{K}}L\mathfrak{g}\mid\operatorname{deg}^{\mathrm{Dr}}_{i}y>0,\ \forall j\neq i\in\mathbb{I}_{0}\colon\operatorname{deg}^{\mathrm{Dr}}_{j}y=0\rangle,$
(8.22) $U_{i-,0}=$ $\ \langle y\in U_{q}^{\mathbb{K}}L\mathfrak{g}\mid\operatorname{deg}^{\mathrm{Dr}}_{i}y<0,\ \forall j\neq i\in\mathbb{I}_{0}\colon\operatorname{deg}^{\mathrm{Dr}}_{j}y=0\rangle.$
Finally, let $U^{i}_{+}$ be the subalgebra of $U_{q}^{\mathbb{K}}L\mathfrak{g}$ generated by $U_{\neq i,+}$ and $U_{i+,0}$. It is the span of all elements which are: either (a) of positive degree along $\alpha_{i}$ and non-negative degree along all the other simple roots, or (b) of arbitrary degree along $\alpha_{i}$ and positive $Q_{i}$-degree. Similarly, let $U^{i}_{-}$ be the subalgebra of $U_{q}^{\mathbb{K}}L\mathfrak{g}$ generated by $U_{\neq i,-}$ and $U_{i-,0}$.
Sometimes we will also consider $U_{q}^{\mathbb{K}}L\mathfrak{g}$ as a $\widehat{Q}$-graded algebra with
$\deg E_{i}=1,\quad\deg F_{i}=-1,\quad\deg K_{i}=\deg\mathbb{K}_{i}=0\qquad(i\in\mathbb{I}),$
where $\widehat{Q}$ is the root lattice of $\widehat{{\mathfrak{g}}}$.
8.2. The braid group action
The Weyl group $W_{0}$ of ${\mathfrak{g}}$ is generated by the simple reflections $s_{i}$ ($i\in\mathbb{I}_{0}$), and acts on $P$ in the usual way. The extended affine Weyl group $\widetilde{W}$ is the semi-direct product $W_{0}\ltimes P$. It contains the affine Weyl group $W=W_{0}\ltimes Q=\langle s_{i}\mid i\in\mathbb{I}\rangle$ as a subgroup, and can also be realized as the semi-direct product
(8.23) $\widetilde{W}\cong\Lambda\ltimes W,$
where $\Lambda=P/Q$ is a finite group of automorphisms of the Dynkin diagram of $\widehat{{\mathfrak{g}}}$. Let $\widetilde{\mathcal{B}}$ be the braid group associated to $\widetilde{W}$.
It is well known [Lus10] that, for each $i\in\mathbb{I}$, there exists an automorphism $T_{i}$ of $U_{q}L\mathfrak{g}$ such that $T_{i}(K_{\mu})=K_{s_{i}\mu}$ and
(8.24) $T_{i}(E_{i})=$ $\ -F_{i}K_{i},\quad$ $T_{i}(E_{j})=$ $\ \sum_{r+s=-a_{ij}}(-1)^{r}q^{-r}E_{i}^{(s)}E_{j}E_{i}^{(r)},$
(8.25) $T_{i}(F_{i})=$ $\ -K_{i}^{-1}E_{i},\quad$ $T_{i}(F_{j})=$ $\ \sum_{r+s=-a_{ij}}(-1)^{r}q^{r}F_{i}^{(r)}F_{j}F_{i}^{(s)},$
for $\mu\in P$ and $i\neq j$. In particular, if $a_{ij}=-1$ then
(8.26) $T_{i}(E_{j})=E_{i}E_{j}-q^{-1}E_{j}E_{i},\qquad T_{i}(F_{j})=F_{j}F_{i}-qF_{i}F_{j}.$
For each $\lambda\in\Lambda$, there is also an automorphism $T_{\lambda}$ such that $T_{\lambda}(E_{i})=E_{\lambda(i)}$, $T_{\lambda}(F_{i})=F_{\lambda(i)}$ and $T_{\lambda}(K_{i})=K_{\lambda(i)}$. These automorphisms satisfy the relations of the braid group $\widetilde{\mathcal{B}}$. We extend this action to an action on $U_{q}^{\mathbb{K}}L\mathfrak{g}$ by setting $T_{i}(\mathbb{K}_{\mu})=\mathbb{K}_{s_{i}\mu}$ and $T_{\lambda}(\mathbb{K}_{i})=\mathbb{K}_{\lambda(i)}$.
8.3. Generalized $q$-Onsager algebras
Baseilhac and Belliard introduced in [BB10] generalized $q$-Onsager algebras, which are coideal subalgebras of $U_{q}\widehat{\mathfrak{g}}$. These algebras are also known in the literature as affine $\imath$quantum groups of split type [LW21]. We first recall the original “Serre-type” presentation of these algebras.
Definition 8.1.
The universal generalized $q$-Onsager algebra $\mathcal{O}_{q}(\widehat{{\mathfrak{g}}})$ is the algebra generated by $B_{i}$ and invertible central elements $\mathfrak{c}_{i}$ $(i\in\mathbb{I})$ subject to relations
[…]
(8.28) $B_{i}^{2}B_{j}-[2]B_{i}B_{j}B_{i}+B_{j}B_{i}^{2}=$ $\ -q\mathfrak{c}_{i}B_{j}\qquad$ if $\mbox{if }a_{ij}=-1,$
[…]
It follows from [Kol14, Theorem 8.3] that $Z(\mathcal{O}_{q}(\widehat{{\mathfrak{g}}}))={\mathbb{C}}[\mathfrak{c}_{i}^{\pm 1}\mid i\in\mathbb{I}]$.
Recently, Lu and Wang [LW21] found a “Drinfeld-type” presentation of $\mathcal{O}_{q}(\widehat{{\mathfrak{g}}})$. We will use the following shorthand notations:
(8.30) $S(r_{1},r_{2}\mid s;i,j)=$ $\ B_{i,r_{1}}B_{i,r_{2}}B_{j,s}-[2]B_{i,r_{1}}B_{j,s}B_{i,r_{2}}+B_{j,s}B_{i,r_{1}}B_{i,r_{2}},$
(8.31) $\mathbb{S}(r_{1},r_{2}\mid s;i,j)=$ $\ S(r_{1},r_{2}\mid s;i,j)+S(r_{2},r_{1}\mid s;i,j),$
(8.32) $R(r_{1},r_{2}\mid s;i,j)=$ $\ q^{2}\mathfrak{c}_{i}\mathfrak{C}^{r_{1}}\big{(}-\sum_{p\geq 0}q^{2p}[2][\Theta_{i,r_{2}-r_{1}-2p-1},B_{j,l-1}]_{q^{-2}}\mathfrak{C}^{p+1}$
(8.33) $\ -\sum_{p\geq 1}q^{2p-1}[2][B_{j,l},\Theta_{i,r_{2}-r_{1}-2p}]_{q^{-2}}\mathfrak{C}^{p}-[B_{j,l},\Theta_{i,r_{2}-r_{1}}]_{q^{-2}},$
(8.34) $\mathbb{R}(r_{1},r_{2}\mid s;i,j)=$ $\ R(r_{1},r_{2}\mid s;i,j)+R(r_{2},r_{1}\mid s;i,j).$
Definition 8.2.
Let ${}^{\mathrm{Dr}}\mathcal{O}_{q}(\widehat{{\mathfrak{g}}})$ be the $\mathbb{C}$-algebra generated by $H_{i,m}$ and $A_{i,r}$, where $m\geq 1$, $r\in{\mathbb{Z}}$, and invertible central elements $\mathfrak{c}_{i}$ ($i\in\mathbb{I}_{0}$), $\mathfrak{C}$, subject to the following relations:
(8.35) $[H_{i,m},H_{j,n}]$ $=0,$
(8.36) $[H_{i,m},A_{j,r}]$ $=\textstyle\frac{[m\cdot a_{ij}]}{m}(A_{j,r+m}-A_{j,r-m}\mathfrak{C}^{m}),$
(8.37) $[A_{i,r},A_{j,s}]$ $=0\quad\mbox{if}\quad a_{ij}=0,$
(8.38) $[A_{i,r},A_{j,s+1}]_{q^{-a_{ij}}}-q^{-a_{ij}}[A_{i,r+1},A_{j,s}]_{q^{a_{ij}}}$ $=0\quad\mbox{if}\quad i\neq j,$
(8.39) $[A_{i,r},A_{i,s+1}]_{q^{-2}}-q^{-2}[A_{i,r+1},A_{i,s}]_{q^{2}}$ $=\mathfrak{c}_{i}\mathfrak{C}^{r}\Theta_{i,s-r+1}-q^{-2}\mathfrak{c}_{i}\mathfrak{C}^{r+1}\Theta_{i,s-r-1}$
$\quad+\mathfrak{c}_{i}\mathfrak{C}^{s}\Theta_{i,r-s+1}-q^{-2}\mathfrak{c}_{i}\mathfrak{C}^{s+1}\Theta_{i,r-s-1},$
(8.40) $\mathbb{S}(r_{1},r_{2}\mid s;i,j)$ $=\mathbb{R}(r_{1},r_{2}\mid s;i,j)\quad\mbox{if}\quad a_{ij}=-1.$
where $m,n\geq 1$; $r,s,r_{1},r_{2}\in{\mathbb{Z}}$; and
$1+\sum_{m=1}^{\infty}(q-q^{-1})\Theta_{i,m}z^{m}=\exp\left((q-q^{-1})\sum_{m=1}^{\infty}H_{i,m}z^{m}\right).$
By convention, $\Theta_{i,0}=(q-q^{-1})^{-1}$ and $\Theta_{i,m}=0$ for $m\leq-1$.
Remark 8.3.
To match our conventions with those of [LW21], set
(8.41) $v=q,\qquad\mathbb{K}_{i}=q^{2}\mathfrak{c}_{i}.$
By [LW21, Theorem 2.16], there is an algebra isomorphism
[…]
This isomorphism and its inverse can be described precisely in terms of a braid group action (see [LW21, §3.3, 3.4] and §8.5 below). We will not need the details in what follows. We note, however, that the isomorphism sends
(8.43) $B_{i}\mapsto A_{i,0},\quad\mathfrak{c}_{i}\mapsto\mathfrak{c}_{i},\quad q^{2N+2}\mathfrak{c}_{\delta}\mapsto\mathfrak{C},$
for $i\in\mathbb{I}_{0}$.
8.4. Coideal structures
The following is a straightforward generalization of §2.5– §2.7 to higher ranks. Given $\mathbf{c}=(c_{0},\cdots,c_{N})\in({\mathbb{C}}^{\times})^{N+1}$, we have a central character
$\chi_{\mathbf{c}}\colon Z(\mathcal{O}_{q}(\widehat{{\mathfrak{g}}}))\to{\mathbb{C}}^{\times},\qquad\mathfrak{c}_{i}\mapsto c_{i}.$
Let $\mathcal{O}_{q}^{\mathbf{c}}(\widehat{{\mathfrak{g}}})$ be the corresponding central reduction, i.e., the quotient of $\mathcal{O}_{q}(\widehat{{\mathfrak{g}}})$ by the ideal generated by $\mathfrak{c}_{i}-c_{i}$.
By [Kol14, Theorem 7.1], for any $\mathbf{s}=(s_{0},\cdots,s_{N})\in{\mathbb{C}}^{N+1}$, there exists an injective algebra homomorphism
(8.44) $\eta_{\mathbf{c},\mathbf{s}}\colon\mathcal{O}_{q}^{\mathbf{c}}(\widehat{{\mathfrak{g}}})\hookrightarrow U_{q}L\mathfrak{g},\quad B_{i}\mapsto e_{i}^{-}-c_{i}e_{i}^{+}K_{i}^{-1}+s_{i}K_{i}^{-1}\quad(i\in\mathbb{I}),$
giving $\mathcal{O}_{q}^{\mathbf{c}}(\widehat{{\mathfrak{g}}})$ the structure of a right coideal subalgebra of $U_{q}L\mathfrak{g}$. There is also a generic version of this embedding
$\eta_{\mathbf{s}}\colon\mathcal{O}_{q}(\widehat{{\mathfrak{g}}})\hookrightarrow U_{q}^{\mathbb{K}}L\mathfrak{g},\quad B_{i}\mapsto e_{i}^{-}-q^{-2}\mathbb{K}_{i}e_{i}^{+}K_{i}^{-1}+s_{i}K_{i}^{-1}\quad(i\in\mathbb{I}).$
If $\mathbf{s}=(0,\cdots,0)$, we abbreviate $\eta=\eta_{\mathbf{s}}$.
We remark that the other results from §2, concerning one-dimensional modules and $C$-duality, as well as the rationality properties from §3 carry over to the $\mathcal{O}_{q}(\widehat{{\mathfrak{g}}})$ case in a natural way.
8.5. The QSP braid group action
By [LW21, Lemma 3.4], for each $i\in\mathbb{I}$, there exists an automorphism $\mathbf{T}_{i}$ of $\mathcal{O}_{q}(\widehat{{\mathfrak{g}}})$ such that $\mathbf{T}_{i}(\mathbb{K}_{\mu})=\mathbb{K}_{s_{i}\mu}$ and
$\mathbf{T}_{i}(B_{j})=\left\{\begin{array}[]{l l}\mathbb{K}_{j}^{-1}B_{j}&\mbox{if }i=j,\\ B_{j}&\mbox{if }a_{ij}=0,\\ B_{j}B_{i}-qB_{i}B_{j}&\mbox{if }a_{ij}=-1,\\ {[2]^{-1}}(B_{j}B_{i}^{2}-q[2]B_{i}B_{j}B_{i}+q^{2}B_{i}^{2}B_{j})+B_{j}\mathbb{K}_{i}&\mbox{if }a_{ij}=-2,\end{array}\right.$
for $\mu\in{\mathbb{Z}}\mathbb{I}$ and $j\in\mathbb{I}$. For each $\lambda\in\Lambda$, there is also an automorphism $\mathbf{T}_{\lambda}$ such that $\mathbf{T}_{\lambda}(B_{i})=B_{\lambda(i)}$ and $\mathbf{T}_{\lambda}(\mathbb{K}_{i})=\mathbb{K}_{\lambda(i)}$. These automorphisms define an action of the braid group $\widetilde{\mathcal{B}}$. We will refer to this action as the Lu–Wang, QSP or coideal subalgebra braid group action to distinguish it from Lusztig’s braid group action from §8.2.
We will need the following lemma.
Lemma 8.4.
We have $\mathbf{T}_{w}(B_{i})=B_{wi}$ for $i\in\mathbb{I}$ and $w\in\widetilde{W}$ such that $wi\in\mathbb{I}$.
Proof.
See [LW21, Lemma 3.5]. ∎
We remark that the analogous statement for the usual braid group action holds as well.
8.6. Rank one subalgebras
Let us recall how the braid group actions can be used to obtain rank one subalgebras in $U_{q}L\mathfrak{g}$ and $\mathcal{O}_{q}(\widehat{{\mathfrak{g}}})$. For $i\in\mathbb{I}_{0}$, let $\omega^{\prime}_{i}=\omega_{i}s_{i}$, and let $U_{i}$ be the subalgebra of $U_{q}L\mathfrak{g}$ generated by
$E_{i},\ F_{i},\ K_{i}^{\pm 1},\ T_{\omega^{\prime}_{i}}(E_{i}),\ T_{\omega^{\prime}_{i}}(F_{i}),\ T_{\omega^{\prime}_{i}}(K_{i}^{\pm 1}).$
By [Bec94, Proposition 3.7], there is an algebra isomorphism $\iota_{i}\colon U_{q}L\mathfrak{sl}_{2}\to U_{i}$ sending
$E_{1}\mapsto E_{i},\ F_{1}\mapsto F_{i},\ K_{1}^{\pm 1}\mapsto K_{i}^{\pm 1},\quad E_{0}\mapsto T_{\omega^{\prime}_{i}}(E_{i}),\ F_{0}\mapsto T_{\omega^{\prime}_{i}}(F_{i}),\ K_{0}^{\pm 1}\mapsto T_{\omega^{\prime}_{i}}(K_{i}^{\pm 1}).$
Let $U_{i}^{\mathbb{K}}$ be the subalgebra of $U_{q}^{\mathbb{K}}L\mathfrak{g}$ generated by $U_{i}$, $\mathbb{K}_{i}^{\pm 1}$ and $(\mathbb{K}_{\delta}\mathbb{K}_{i}^{-1})^{\pm 1}$. The algebra isomorphism $\iota_{i}$ extends to an isomorphism $\iota_{i}\colon U_{q}^{\mathbb{K}}L\mathfrak{sl}_{2}\to U_{i}^{\mathbb{K}}$ sending $\mathbb{K}_{1}\mapsto K_{i}$ and $\mathbb{K}_{0}\mapsto\mathbb{K}_{\delta}\mathbb{K}_{i}^{-1}$. For future use, we note that it follows easily from the definitions that
(8.45) $\iota_{i}(x^{-}_{r})=x^{-}_{i,r}.$
Moreover, for $i\in\mathbb{I}_{0}$, let $\mathcal{O}_{q}^{[i]}$ be the subalgebra of $\mathcal{O}_{q}(\widehat{{\mathfrak{g}}})$ generated by $B_{i}$, $\mathbf{T}_{\omega^{\prime}_{i}}(B_{i})$, $\mathbb{K}_{i}^{\pm 1}$ and $(\mathbb{K}_{\delta}\mathbb{K}_{i}^{-1})^{\pm 1}$. By [LW21, Proposition 3.9], there is an algebra isomorphism $\iota\colon\mathcal{O}_{q}\to\mathcal{O}_{q}^{[i]}$ sending
(8.46) $B_{1}\mapsto B_{i},\quad B_{0}\mapsto\mathbf{T}_{\omega^{\prime}_{i}}(B_{i}),\quad\mathbb{K}_{1}\mapsto\mathbb{K}_{i},\quad\mathbb{K}_{0}\mapsto\mathbb{K}_{\delta}\mathbb{K}_{i}^{-1}.$
We also note that $A_{-1}$ is sent to $\mathbf{T}_{\omega_{i}}(B_{i})$.
9. Compatibility of braid group actions
The braid group action on $\mathcal{O}_{q}(\widehat{{\mathfrak{g}}})$ is not compatible with the action on $U_{q}L\mathfrak{g}$, in the sense that the former is not the restriction of the latter. Similarly, the rank one subalgebras from §8.6 are not compatible, i.e., for each $i\in\mathbb{I}_{0}$, the diagram
[…]
does not commute. Nevertheless, the goal of this section is to show that, in an appropriate sense, diagram (9.1) does commute modulo the “Drinfeld positive half” of $U_{q}^{\mathbb{K}}L\mathfrak{g}$.
9.1. Action on generators
Assume that ${\mathfrak{g}}$ is not of type $\mathsf{A}_{1}$. Set
$\tilde{E}_{i}=-q^{-2}\mathbb{K}_{i}E_{i}K_{i}^{-1}.$
We begin by investigating the image of elements such as $\mathbf{T}_{i}(B_{i\pm 1})$ inside $U_{q}^{\mathbb{K}}L\mathfrak{g}$.
Lemma 9.1.
If $a_{ij}=-1$ then
$\eta(\mathbf{T}_{i}(B_{j}))=T_{i}(\eta(B_{j}))+(q-q^{-1})F_{i}\tilde{E}_{j}=T_{i}(\eta(B_{j}))+[\tilde{E}_{j},F_{i}]_{q}.$
Proof.
If $a_{ij}=-1$ then
(9.2) $T_{i}(\eta(B_{j}))=(F_{j}F_{i}-qF_{i}F_{j})-q^{-2}\mathbb{K}_{i}\mathbb{K}_{j}(E_{i}E_{j}-q^{-1}E_{j}E_{i})K_{i}^{-1}K_{j}^{-1}$
and
(9.3) $\eta(\mathbf{T}_{i}(B_{j}))=$ $\ \eta(B_{j}B_{i}-qB_{i}B_{j})$
(9.4) $=$ $\ (F_{j}-q^{-2}\mathbb{K}_{j}E_{j}K_{j}^{-1})(F_{i}-q^{-2}\mathbb{K}_{i}E_{i}K_{i}^{-1})$
(9.5) $\ -q(F_{i}-q^{-2}\mathbb{K}_{i}E_{i}K_{i}^{-1})(F_{j}-q^{-2}\mathbb{K}_{j}E_{j}K_{j}^{-1})$
(9.6) $=$ $\ (F_{j}F_{i}-qF_{i}F_{j})-q^{-2}\mathbb{K}_{i}\mathbb{K}_{j}(E_{i}E_{j}-q^{-1}E_{j}E_{i})K_{i}^{-1}K_{j}^{-1}$
(9.7) $\ +(q-q^{-1})q^{-2}\mathbb{K}_{j}F_{i}E_{j}K_{j}^{-1},$
since
(9.8) $F_{j}E_{i}K_{i}^{-1}-qE_{i}K_{i}^{-1}F_{j}=0.$
Applying (8.26) concludes the proof. ∎
For future use, let us record (9.8) as
(9.9) $[F_{i},\tilde{E}_{i\pm 1}]=0.$
9.2. Type $\mathsf{A}$ combinatorics
From now on assume that ${\mathfrak{g}}$ is of type $\mathsf{A}_{N}$ with $N\geq 2$. Then $\Lambda=\langle\pi\rangle$, where $\pi$ is the automorphism of order $N+1$ of the affine Dynkin diagram of type $\mathsf{A}_{N}^{(1)}$ sending $i\mapsto i+1\mod N+1$. The highest root is $\theta=\alpha_{1}+\alpha_{2}+\cdots+\alpha_{n}$.
Let $w_{0}$ be the longest element of the finite Weyl group $W_{0}$ of ${\mathfrak{g}}$. Given $i\in\mathbb{I}_{0}$, let $w_{i}$ the longest element of the Weyl group corresponding to the Dynkin diagram of ${\mathfrak{g}}$ with the node labelled by $i$ removed. Given $1\leq k<l$, let $[k,l]=s_{k}s_{k+1}\cdots s_{l}$.
Proposition 9.2.
Let $\widetilde{W}$ be the extended affine Weyl group of type $\mathsf{A}_{N}^{(1)}$, for $N\geq 1$. Then
$\omega_{i}=\pi^{i}[N-i+1,N]\cdots[2,i+1][1,i].$
This is a reduced expression.
Proof.
It follows from [Bou02, Ch. VI §2.3, Theorem 6] that
$\omega_{i}=\pi^{i}w_{0}w_{i}.$
It is easy to check that $w_{0}w_{i}=\tau^{N+1-i}$, where $\tau$ is the inverse of the cyclic permutation $(12\cdots N+1)$, if we identify $W$ with the symmetric group $S_{N+1}$. It is well known that such permutations have an expression of the form $[N-i+1,N]\cdots[2,i+1][1,i]$. To show that this expression is reduced, it suffices to check that $\ell(\omega_{i})=i(N-i+1)$. By [IM65, Proposition 1.23], $\ell(\omega_{i})$ is equal to the number of positive roots which contain $\alpha_{i}$ in their expression in terms of the basis of simple roots. Such roots are precisely $\alpha_{k}+\cdots+\alpha_{l}$, for $1\leq k\leq i$ and $i\leq l\leq N$, completing the proof. ∎
For example, let $N=5$. Then:
$\omega_{1}\mapsto\pi s_{5}s_{4}s_{3}s_{2}s_{1},\quad\omega_{2}\mapsto\pi^{2}s_{4}s_{5}s_{3}s_{4}s_{2}s_{3}s_{1}s_{2},\quad\omega_{3}\mapsto\pi^{3}s_{3}s_{4}s_{5}s_{2}s_{3}s_{4}s_{1}s_{2}s_{3},$
$\omega_{4}\mapsto\pi^{4}s_{2}s_{3}s_{4}s_{5}s_{1}s_{2}s_{3}s_{4},\quad\omega_{5}\mapsto\pi^{5}s_{1}s_{2}s_{3}s_{4}s_{5}.$
9.3. Inductive arguments
Define non-commutative polynomials $P_{k}(y_{1},\cdots,y_{k})$ over ${\mathbb{C}}$ by induction in the following way:
(9.10) $P_{1}(y_{1})=$ $\ y_{1},\qquad$ $P_{k+1}(y_{1},\cdots,y_{k})=$ $\ [P_{k}(y_{1},\cdots,y_{k}),y_{k+1}]_{q}.$
Clearly, we have
(9.11) $P_{k}(y_{1},\cdots,y_{k})=$ $\ P_{k-l+1}(P_{l}(y_{1},\cdots,y_{l}),y_{l+1},\cdots,y_{k}).$
for any $11\leq l\leq k-$ 1. We will need the following lemma.
Lemma 9.3.
Let $2\leq m\leq N$. Then $\mathbf{T}_{m}\cdots\mathbf{T}_{2}(B_{1})=P_{m}(B_{1},\cdots,B_{m})$. Moreover,
$\eta(\mathbf{T}_{m}\cdots\mathbf{T}_{2}(B_{1}))=\sum_{\begin{subarray}{c}\varepsilon_{k}\geq\varepsilon_{k+1},\\ k=1,...,m-1\end{subarray}}P_{m}(\tilde{e}_{1}^{\varepsilon_{1}},\cdots,\tilde{e}_{m}^{\varepsilon_{m}}),$
where each $\varepsilon_{l}\in\{+,-\}$, and $\tilde{e}_{i}^{+}=\tilde{E}_{i}$, $\tilde{e}_{i}^{-}=F_{i}$. In other words, the RHS consists of polynomials in letters $\tilde{E}_{1},\cdots,\tilde{E}_{l},F_{l+1},\cdots,F_{m}$ (with each letter occurring exactly once), for $0\leq l\leq m$. In particular, these polynomials have degree $\sum_{j=1}^{l}\alpha_{j}-\sum_{j=l+1}^{m}\alpha_{j}$.
Proof.
We prove the statement by induction on $m$. The case $m=2$ reduces to Lemma 9.1. Otherwise, by induction,
$\mathbf{T}_{m-1}\cdots\mathbf{T}_{2}(B_{1})=P_{m-1}(B_{1},\cdots,B_{m-1}).$
Then
(9.12) $\mathbf{T}_{m}\cdots\mathbf{T}_{2}(B_{1})=$ $\ P_{m-1}(B_{1},\cdots,B_{m-2},\mathbf{T}_{m}(B_{m-1}))$
(9.13) $=$ $\ [P_{m-1}(B_{1},\cdots,B_{m-1}),B_{m}]_{q}$
(9.14) $=$ $\ P_{m}(B_{1},\cdots,B_{m}),$
since $B_{m}$ commutes with $B_{1},\cdots,B_{m-2}$. This proves the first statement of the lemma.
Moreover, by induction,
$\eta(P_{m-1}(B_{1},\cdots,B_{m-1}))=\sum_{\varepsilon_{k}\geq\varepsilon_{k+1}}P_{m-1}(\tilde{e}_{1}^{\varepsilon_{1}},\cdots,\tilde{e}_{{m-1}}^{\varepsilon_{m-1}}).$
Therefore, it suffices to show that
$P_{m}(\tilde{e}_{1}^{\varepsilon_{1}},\cdots,\tilde{e}_{{m-2}}^{\varepsilon_{m-2}},F_{m-1},\tilde{E}_{m})=0.$
But this is indeed the case since
$P_{m}(\tilde{e}_{1}^{\varepsilon_{1}},\cdots,\tilde{e}_{{m-2}}^{\varepsilon_{m-2}},F_{m-1},\tilde{E}_{m})=P_{m-1}(\tilde{e}_{1}^{\varepsilon_{1}},\cdots,\tilde{e}_{{m-2}}^{\varepsilon_{m-2}},[F_{m-1},\tilde{E}_{m}]_{q})$
and $[F_{m-1},\tilde{E}_{m}]_{q}=0$ by (9.9). ∎
Lemma 9.3 has the following “dual”.
Lemma 9.4.
Let $2\leq m\leq N$. Then $\mathbf{T}_{1}\cdots\mathbf{T}_{m-1}(B_{m})=P_{m}(B_{m},\cdots,B_{1})$. Moreover,
$\eta(\mathbf{T}_{1}\cdots\mathbf{T}_{m-1}(B_{m}))=\sum_{\begin{subarray}{c}\varepsilon_{k}\leq\varepsilon_{k+1},\\ k=1,...,m-1\end{subarray}}P_{m}(\tilde{e}_{m}^{\varepsilon_{m}},\cdots,\tilde{e}_{1}^{\varepsilon_{1}}),$
In other words, the RHS can be expressed as a sum of polynomials consisting of letters $\tilde{E}_{m},\cdots,\tilde{E}_{l},F_{l-1},\cdots,F_{1}$, for $11\leq l\leq m+$ 1. In particular, these polynomials have degree $-\sum_{j=1}^{l-1}\alpha_{j}+\sum_{j=l}^{m}\alpha_{j}$.
Proof.
The proof is analogous to that of Lemma 9.3. ∎
9.4. Weak compatibility
We first recall the following fact.
Lemma 9.5.
We have $\operatorname{deg}^{\mathrm{Dr}}E_{0}=-\theta=-\sum_{j=1}^{N}\alpha_{j}$ and $\operatorname{deg}^{\mathrm{Dr}}F_{0}=\theta=\sum_{j=1}^{N}\alpha_{j}$.
Proof.
This follows directly from the explicit description of the images of $E_{0}$ and $F_{0}$ in the Drinfeld realization, see, e.g., [Bec94, Remark 4.7] and [CP94, Theorem 2.2]. ∎
Let us write $\eta(\mathbf{T}_{\omega^{\prime}_{i}}(B_{i}))$ as a sum of $Q$-homogeneous terms. We say that a term is irrelevant if its degree is not ±$\pm\alpha_{i}$. The proposition below is our first result on the compatibility of the two braid group actions.
Proposition 9.6.
Let $1\leq i\leq N$. Every irrelevant term of $\eta(\mathbf{T}_{\omega^{\prime}_{i}}(B_{i}))$ lies in $U_{d_{i}=1,+}$. Hence
(9.15) $\eta(\mathbf{T}_{\omega_{i}^{\prime}}(B_{i}))$$\equiv$ $\ T_{\omega_{i}^{\prime}}(\eta(B_{i}))$ $\mod U_{d_{i}=1,+}.$
Proof.
By Lemma 9.4,
$\mathbf{T}_{1}\cdots\mathbf{T}_{i-1}(B_{i})=[P_{i-1}(B_{i},\cdots,B_{2}),B_{1}]_{q}.$
Next, observe that
(9.16) $(s_{N-i+1}\cdots s_{N})\cdots(s_{2}\cdots s_{i+1})(\alpha_{1})=$ $\ s_{N-i+1}\cdots s_{2}(\alpha_{1}),$
(9.17) $(s_{N-i+1}\cdots s_{N})\cdots(s_{2}\cdots s_{i+1})(\alpha_{k})=$ $\ \alpha_{k+N-i},$
for $2\leq k\leq B_{i}$. Therefore, Lemmas 8.4 and 9.3-9.4 imply that
(9.18) $\mathbf{T}_{\omega^{\prime}_{i}}(B_{i})=$ $\ \pi^{i}[P_{i-1}(B_{N},\cdots,B_{N-i+2}),\mathbf{T}_{N-i+1}\cdots\mathbf{T}_{2}(B_{1})]_{q}$
(9.19) $=$ $\ \pi^{i}[P_{i-1}(B_{N},\cdots,B_{N-i+2}),P_{N-i+1}(B_{1},\cdots,B_{N-i+1})]_{q}$
(9.20) $=$ $\ \pi^{i}\sum_{\begin{subarray}{c}\varepsilon_{k}\leq\varepsilon_{k+1},\\ k=N-i+2,...,N\end{subarray}}\sum_{\begin{subarray}{c}\varepsilon_{l}\geq\varepsilon_{l+1},\\ l=1,...,N-i\end{subarray}}[P_{i-1}(\tilde{e}_{N}^{\varepsilon_{N}},\cdots,\tilde{e}_{N-i+2}^{\varepsilon_{N-i+2}}),P_{N-i+1}(\tilde{e}_{1}^{\varepsilon_{1}},\cdots,\tilde{e}_{N-i+1}^{\varepsilon_{N-i+1}})]_{q}$
(9.21) $=$ $\ \sum_{\begin{subarray}{c}\varepsilon_{k}\leq\varepsilon_{k+1},\\ k=1,...,i-1\end{subarray}}\sum_{\begin{subarray}{c}\varepsilon_{l}\geq\varepsilon_{l+1}\geq\varepsilon_{0},\\ l=i+1,...,N-1\end{subarray}}[P_{i-1}(\tilde{e}_{i-1}^{\varepsilon_{i-1}},\cdots,\tilde{e}_{1}^{\varepsilon_{1}}),P_{N-i+1}(\tilde{e}_{i+1}^{\varepsilon_{i+1}},\cdots,\tilde{e}_{0}^{\varepsilon_{0}})]_{q}.$
By Lemma 9.5, each irrelevant term inside of the summation in (9.21) is already in $U_{d_{i}=1,+}$ unless $\varepsilon_{0}=+$. So suppose that $\varepsilon_{0}=+$ and that at least one of the superscripts $\varepsilon_{k}=-$. This forces $\varepsilon_{1}=-$ and $\varepsilon_{i+1}=\cdots=\varepsilon_{0}=+$. We then have
(9.23) $[P_{i-1}(\tilde{e}_{i-1}^{\varepsilon_{i-1}},\cdots,\tilde{e}_{1}^{\varepsilon_{1}}),$ $\ P_{N-i+1}(\tilde{e}_{i+1}^{\varepsilon_{i+1}},\cdots,\tilde{e}_{0}^{\varepsilon_{0}})]_{q}=$
(9.24) $=$ $\ [[P_{i-2}(\tilde{e}_{i-1}^{\varepsilon_{i-1}},\cdots,\tilde{e}_{2}^{\varepsilon_{2}}),F_{1}]_{q},[P_{N-i}(\tilde{E}_{i+1},\cdots,\tilde{E}_{N}),\tilde{E}_{0}]_{q}]_{q}$
(9.25) $=$ $\ [P_{N-i}(\tilde{E}_{i+1},\cdots,\tilde{E}_{N}),[P_{i-2}(\tilde{e}_{i-1}^{\varepsilon_{i-1}},\cdots,\tilde{e}_{2}^{\varepsilon_{2}}),[F_{1},\tilde{E}_{0}]_{q}]_{q}]_{q}$
[…]
since, by (9.9), $[F_{1},\tilde{E}_{0}]_{q}=0$. We conclude that every irrelevant term in (9.21) is indeed in $U_{d_{i}=1,+}$. Moreover, the relevant terms are precisely
(9.27) $[P_{i-1}(\tilde{E}_{i-1},\cdots,\tilde{E}_{1}),P_{N-i+1}(\tilde{E}_{i+1},\cdots,\tilde{E}_{0})]_{q}=$ $\ T_{\omega_{i}^{\prime}}(\tilde{E}_{i}),$
(9.28) $[P_{i-1}({F}_{i-1},\cdots,{F}_{1}),P_{N-i+1}({F}_{i+1},\cdots,{F}_{0})]_{q}=$ $\ T_{\omega_{i}^{\prime}}({F}_{i}).$
This concludes the proof. ∎
Corollary 9.7 (Compatibility theorem).
The diagram (9.1) commutes modulo $U_{\neq i,+}$, for each $1\leq i\leq N$.
Proof.
Let $\mu$ denote the composition of the east and south arrows, and $\nu$ the composition of the south and east arrows in (9.1). We have $\mu(B_{0})=\nu(B_{0})$ and, by Proposition 9.6, $\mu(B_{-1})\equiv\nu(B_{-1})$ modulo $U_{d_{i}=1,+}$. Since $B_{0}$ and $B_{-1}$ generate $\mathcal{O}_{q}$, and $\mu(B_{0})$ contains terms of degree ±$\pm\alpha_{i}$, it follows that the diagram commutes modulo $U_{\neq i,+}$. ∎
9.5. Strong compatibility
The goal of this section is to strengthen the results of the preceding subsection to pave the way for a higher rank generalization of the factorization theorem. Set $\mathfrak{C}_{i}=\mathfrak{C}^{-1}\mathbb{K}_{i}$. It is easily verified that
$\mathbf{T}_{\omega^{\prime}_{i}}(\mathbb{K}_{i})=\mathfrak{C}_{i}^{-1},\qquad T_{\omega^{\prime}_{i}}(K_{i})=K_{i}^{-1}.$
In the calculations below, we will frequently make use of a remainder term $Q_{i}$. The statement of the following lemma contains its definition.
Lemma 9.8.
For each $i\in\mathbb{I}_{0}$:
$\eta(A_{i,-1})=x^{-}_{i,1}-\mathfrak{C}_{i}x^{+}_{i,-1}K_{i}+Q_{i},\qquad Q_{i}\in U_{d_{i}\geq 1,+}.$
Hence
$\eta(A_{i,-1})\equiv\iota_{i}\eta(A_{-1})\quad\mod U_{d_{i}\geq 1,+}.$
Proof.
By [Bec94, Proposition 4.7] [^4] More precisely, in the first equality below, an adjustment by a sign function $o(i)$ which alternates on adjacent vertices is necessary. An analogous adjustment occurs in the formula for $A_{i,-1}$. Since we work with a fixed vertex $i\in\mathbb{I}_{0}$ most of the time, we may assume for simplicity that $o(i)=1$.,
(9.29) $x_{i,-1}^{+}=$ $\ T_{\omega_{i}}(E_{i})=T_{\omega^{\prime}_{i}}(-F_{i}K_{i})=-T_{\omega^{\prime}_{i}}(F_{i})K_{i}^{-1},$
(9.30) $x_{i,1}^{-}=$ $\ T_{\omega_{i}}(F_{i})=T_{\omega^{\prime}_{i}}(-K_{i}^{-1}E_{i})=-K_{i}T_{\omega^{\prime}_{i}}(E_{i}).$
Moreover, by [LW21, Theorem 3.13],
$A_{i,-1}=\mathbf{T}_{\omega_{i}}(B_{i})=\mathbf{T}_{\omega^{\prime}_{i}}(\mathbb{K}_{i}^{-1}B_{i})=\mathfrak{C}_{i}\mathbf{T}_{\omega^{\prime}_{i}}(B_{i}).$
Hence, by Proposition 9.6 and (9.29)–(9.30),
$\eta(A_{i,-1})=\mathfrak{C}_{i}T_{\omega^{\prime}_{i}}(F_{i}-K_{i}^{-1}\mathbb{K}_{i}E_{i})+Q_{i}=-\mathfrak{C}_{i}x^{+}_{i,-1}K_{i}+x^{-}_{i,1}+Q_{i},$
completing the proof. ∎
Lemma 9.9.
We have
$\eta(H_{i,1})\equiv\iota_{i}\eta(H_{1})+q^{2}\mathfrak{C}_{i}^{-1}[Q_{i},F_{i}]_{q^{-2}}\quad\mod U_{d_{i}\geq 2,+}.$
Proof.
We have
(9.31) $\eta(H_{i,1})=$ $\ q^{2}\mathfrak{C}_{i}^{-1}\eta([A_{i,-1},A_{i,0}]_{q_{i}^{-2}})$
(9.32) $=$ $\ q^{2}\mathfrak{C}_{i}^{-1}[\iota_{i}\eta(A_{-1})+Q_{i},\iota_{i}\eta(A_{0})]_{q^{-2}}$
(9.33) $=$ $\ \iota_{i}\eta(H_{1})+q^{2}\mathfrak{C}_{i}^{-1}[Q_{i},F_{i}+\tilde{E}_{i}]_{q^{-2}}.$
Since $[Q_{i},\tilde{E}_{i}]_{q^{-2}}\in U_{d_{i}\geq 2,+}$, the lemma follows. ∎
Proposition 9.10.
For all $r\in{\mathbb{Z}}$:
$\eta(A_{i,r})\equiv\iota_{i}\eta(A_{r})\quad\mod U_{d_{i}\geq 1,+}.$
Proof.
The result holds for $r=-1,0$ by Proposition 9.6. The general case is proven using upward and downward induction. We explicitly do the former, the latter being analogous. Relation (8.36) implies that
(9.34) $A_{i,r+1}=[2]_{q}^{-1}[H_{i,1},A_{i,r}]+\mathfrak{C}A_{i,r-1}.$
Using Lemma 9.9, let us write $\eta(H_{i,1})=\iota_{i}\eta(H_{1})+q^{2}\mathfrak{C}_{i}^{-1}[Q_{i},F_{i}]_{q^{-2}}+R$, with $R\in U_{d_{i}\geq 2,+}$. By induction, we may assume that
$\eta(A_{i,r-1})\equiv\iota_{i}\eta(A_{r-1}),\quad[\iota_{i}\eta(H_{1}),\eta(A_{i,r})]\equiv\iota_{i}\eta([H_{1},A_{r}])$
modulo $U_{d_{i}\geq 1,+}$. Therefore, it suffices to show that
$[q^{2}\mathfrak{C}_{i}^{-1}[Q_{i},F_{i}]_{q^{-2}}+R,\eta(A_{i,r})]\in U_{d_{i}\geq 1,+}.$
By induction, we can write $\eta(A_{i,r})=\iota_{i}\eta(A_{r})+S$, with $S\in U_{d_{i}\geq 1,+}$. On the other hand, by Lemma 4.5 and (8.45), $\iota_{i}\eta(A_{r})=x_{i,-r}^{-}+S^{\prime}$ with $S^{\prime}\in U_{[i],+}$, where $U_{[i],+}$ is the subalgebra of $U_{[i]}$ spanned by elements of positive degree (with respect to $\alpha_{i}$). We observe immediately that
$[R,\eta(A_{i,r})]+[q^{2}\mathfrak{C}_{i}^{-1}[Q_{i},F_{i}]_{q^{-2}},S+S^{\prime}]\in U_{d_{i}\geq 1,+}.$
Therefore, it suffices to show that
(9.35) $[[Q_{i},F_{i}]_{q^{-2}},x_{i,-r}^{-}]\in U_{d_{i}\geq 1,+}.$
In fact, we show that (9.35) vanishes. This will be achieved in §9.6 below. ∎
9.6. Completion of the proof
We first state some auxiliary lemmas. Define the non-commutative polynomial $P^{\prime}_{k}(y_{1},\cdots,y_{k})$ over ${\mathbb{C}}$ by induction in the following way:
(9.36) $P^{\prime}_{1}(y_{1})=$ $\ y_{1},\qquad$ $P^{\prime}_{k+1}(y_{1},\cdots,y_{k})=$ $\ P^{\prime}_{k}(y_{1},\cdots,y_{k-1},[y_{k},y_{k+1}]_{q}),$
We have
(9.37) $P^{\prime}_{k}(y_{1},\cdots,y_{k})=$ $\ P^{\prime}_{l+1}(y_{1},\cdots,y_{l},P^{\prime}_{k-l}(y_{l+1},\cdots,y_{k})),$
for any $11\leq l\leq k-$ 1.
We say that a tuple $(y_{1},\cdots,y_{k})$ is almost commuting if $y_{m}y_{n}=y_{n}y_{m}$, for $|m-n|>1$.
Lemma 9.11.
The polynomials $P_{k}$ and $P^{\prime}_{k}$ are equal if $(y_{1},\cdots,y_{k})$ is almost commuting.
Proof.
The first statement is proven by induction. By definition, $P_{1}=P^{\prime}_{1}$ and $P_{2}=P^{\prime}_{2}$. Moreover,
(9.39) $P^{\prime}_{k+1}(y_{1},\cdots,y_{k+1})=$ $\ P^{\prime}_{2}(y_{1},P^{\prime}_{k}(y_{2},\cdots,y_{k+1}))$
(9.40) $=$ $\ P^{\prime}_{2}(y_{1},P_{k}(y_{2},\cdots,y_{k+1}))$
(9.41) $=$ $\ P_{k}(P^{\prime}_{2}(y_{1},y_{2}),y_{3},\cdots,y_{k+1})$
(9.42) $=$ $\ P_{k}(P_{2}(y_{1},y_{2}),y_{3},\cdots,y_{k+1})=P^{\prime}_{k+1}(y_{1},\cdots,y_{k+1}).\qed$
Lemma 9.12.
Let $12\leq i\leq n-$ 1. Then
  1. (1)
    $[\tilde{E}_{i},[\tilde{E}_{i-1},Y]_{q}]_{q}=q[E_{i},[\tilde{E}_{i-1}K_{i}^{-1},Y]_{q}]_{q^{2}}$ if $[K_{i}^{-1},Y]=0$,
  2. (2)
    $[[\tilde{E}_{i-1},Y]_{q},F_{i}]_{q^{-2}}=[\tilde{E}_{i-1},[Y,F_{i}]_{q^{-1}}]$,
  3. (3)
    $[[\tilde{E}_{i+1},Y]_{q},F_{i}]_{q^{-1}}=[\tilde{E}_{i+1},[Y,F_{i}]]$,
for any $Y\in U_{q}^{\mathbb{K}}L\mathfrak{g}$. If $[Y,F_{i}]=0$ then
(9.43) $[[\tilde{E}_{i-1},[\tilde{E}_{i+1},Y]_{q}]_{q},F_{i}]_{q^{-2}}=0.$
Proof.
The first three statements follow by a direct calculation. The fourth statement follows from the second and third if we substitute $[\tilde{E}_{i+1},Y]_{q}$ for the $Y$ in (2). ∎
Lemma 9.13.
Let $i-1\leq k\leq 1$. The element
$Z=[[P_{i-1}(\tilde{E}_{i-1},\cdots,\tilde{E}_{k},F_{k-1},\cdots,F_{1}),P_{N-i+1}(F_{i+1},\cdots,F_{N},F_{0})]_{q},F_{i}]_{q^{-2}}$
can be expressed as a polynomial in $x_{j,r}^{+}$ $(j\neq i)$, with coefficients in ${\mathbb{C}}$. Hence
(9.44) $[Z,x_{i,s}^{-}]=0$
for all $s\in{\mathbb{Z}}$.
Similarly, if $i+1\leq k\leq n$ or $k=0$, then
$Z^{\prime}=[[P_{i-1}(F_{i-1},\cdots,F_{1}),P_{N-i+1}(\tilde{E}_{i+1},\cdots,\tilde{E}_{k},F_{k-1},\cdots F_{N},F_{0})]_{q},F_{i}]_{q^{-2}}$
can also be expressed as a polynomial in $x_{j,r}^{+}$ $(j\neq i)$, with coefficients in ${\mathbb{C}}$. Hence
(9.45) $[Z^{\prime},x_{i,s}^{-}]=0.$
Proof.
We will only prove the statement for $Z$, the one for $Z^{\prime}$ being analogous. The equality (9.44) follows directly from relation (8.13) and $Z$ being a polynomial in $x_{j,r}^{+}$ $(j\neq i)$. Therefore, it suffices to prove the latter. Using (9.37) and Lemma 9.11, we have
(9.46) $[P_{i-1}(\tilde{E}_{i-1},\cdots,\tilde{E}_{k},$ $\ F_{k-1},\cdots,F_{1}),P_{N-i+1}(F_{i+1},\cdots,F_{N},F_{0})]_{q}=$
(9.47) $=$ $\ P^{\prime}_{i}(\tilde{E}_{i-1},\cdots,\tilde{E}_{k},F_{k-1},\cdots,F_{1},P^{\prime}_{N-i+1}(F_{i+1},\cdots,F_{N},F_{0}))$
(9.48) $=$ $\ P^{\prime}_{N}(\tilde{E}_{i-1},\cdots,\tilde{E}_{k},F_{k-1},\cdots,F_{1},F_{i+1},\cdots,F_{N},F_{0}).$
In the following, let us abbreviate
$\vec{F}$ ” $=$$F_{k{\shortminus}1},\cdots,F_{1},F_{i},F_{i+1},\cdots,F_{N},F_{0}$”.
Direct calculation using Lemma 9.12 shows that
(9.49) $Z=$ $\ {\shortminus}q[\tilde{E}_{i{\shortminus}1},P^{\prime}_{N}(F_{i},\tilde{E}_{i{\shortminus}2},\cdots,\tilde{E}_{k},F_{k{\shortminus}1},\cdots,F_{1},F_{i+1},\cdots,F_{N},F_{0})]$
(9.50) $=$ $\ {\shortminus}q[\tilde{E}_{i{\shortminus}1},P^{\prime}_{N}(\tilde{E}_{i{\shortminus}2},\cdots,\tilde{E}_{k},F_{k{\shortminus}1},\cdots,F_{1},F_{i},F_{i+1},\cdots,F_{N},F_{0})]$
(9.51) $=$ $\ {\shortminus}q[\tilde{E}_{i{\shortminus}1},[E_{i{\shortminus}2},P^{\prime}_{N{\shortminus}1}(E_{i{\shortminus}3},\cdots,E_{k}K_{i{\shortminus}2}^{-1}\cdots K_{k}^{-1},\vec{F})]]$
(9.52) $=$ $\ {\shortminus}q^{2}[E_{i{\shortminus}1},[E_{i{\shortminus}2}K_{i{\shortminus}1}^{-1},P^{\prime}_{N{\shortminus}1}(E_{i{\shortminus}3},\cdots,E_{k}K_{i{\shortminus}2}^{-1}\cdots K_{k}^{-1},\vec{F})]_{q^{-1}}]_{q^{2}}$
(9.53) $=$ $\ {\shortminus}q^{2}[E_{i{\shortminus}1},[E_{i{\shortminus}2},[E_{i{\shortminus}3}K_{i{\shortminus}1}^{-1},P^{\prime}_{N{\shortminus}2}(E_{i{\shortminus}4},\cdots,E_{k}K_{i{\shortminus}2}^{-1}\cdots K_{k}^{-1},\vec{F})]]_{q^{-1}}]_{q^{2}}$
(9.54) $=$ $\ {\shortminus}q^{2}[E_{i{\shortminus}1},[E_{i{\shortminus}2},P^{\prime}_{N{\shortminus}1}(E_{i{\shortminus}3},E_{i{\shortminus}4},\cdots,E_{k}K_{i{\shortminus}1}^{-1}\cdots K_{k}^{-1},\vec{F})]_{q^{-1}}]_{q^{2}}$
(9.55) $=$ $\ q^{3}(q{\shortminus}q^{-1})[E_{i{\shortminus}1},[E_{i{\shortminus}2},P^{\prime}_{i{\shortminus}k{\shortminus}1}(E_{i{\shortminus}3},E_{i{\shortminus}4},\cdots,$
(9.56) $\ E_{k+1},E_{k}P^{\prime}_{N{\shortminus}i+k+2}(\vec{F})K_{i{\shortminus}1}^{-1}\cdots K_{k}^{-1})]_{q^{-1}}]_{q^{2}}.$
Therefore, it suffices to show that $P^{\prime}_{N{\shortminus}i+k+2}(\vec{F})K_{i{\shortminus}1}^{-1}\cdots K_{k}^{-1}$ can be expressed as a polynomial in $x_{j,r}^{+}$ $(j\neq i)$. First assume that $i>1$. We claim that
$P^{\prime}_{N{\shortminus}i+k+2}(\vec{F})K_{i{\shortminus}1}^{-1}\cdots K_{k}^{-1}=P^{\prime}_{i-k}(x_{i-1,-1}^{+},E_{i-2},\cdots,E_{k})=:Y.$
Let
$\zeta=\pi^{i-1}[N-i+2,N]\cdots[2,i].$
Indeed,
(9.57) $Y=$ $\ P^{\prime}_{i-k}(x_{i-1,-1}^{+},E_{i-2},\cdots,E_{k})=T_{\omega_{i-1}}P^{\prime}_{i-k}(E_{i-1},E_{i-2},\cdots,E_{k})$
(9.58) $=$ $\ T_{\omega_{i-1}}T_{k}\cdots T_{i-2}(E_{i-1})=T_{\zeta}T_{[1,k]}(E_{k})$
(9.59) $=$ $\ -T_{\zeta}T_{[1,k-1]}(F_{k}K_{k})=-T_{\zeta}(P^{\prime}_{k}(F_{k},\cdots,F_{1})K_{1}\cdots K_{k})$
(9.60) $=$ $\ -T_{\pi^{i-1}}(P^{\prime}_{k}(F_{N+k-i+1},\cdots,F_{N-i+3},F_{1},\cdots,F_{N-i+2})\cdot$
(9.61) $\ \cdot K_{N+k-i+1}\cdots K_{N-i+3}K_{1}\cdots K_{N-i+2})$
(9.62) $=$ $\ -P^{\prime}_{k}(F_{k-1},\cdots,F_{1},F_{i},\cdots,F_{N},F_{0})K_{k-1}\cdots K_{1}K_{i}\cdots K_{N}K_{0}$
(9.63) $=$ $\ -P^{\prime}_{N{\shortminus}i+k+2}(\vec{F})K_{i{\shortminus}1}^{-1}\cdots K_{k}^{-1},$
proving the claim. The case $i=1$ is similar but easier, so we leave it to the reader. ∎
Proposition 9.14.
We have
$[[Q_{i},F_{i}]_{q^{-2}},x^{-}_{i,r}]=0.$
Proof.
By the definition of $Q_{i}$ and (9.21), we have
(9.64) $Q_{i}=\mathfrak{C}_{i}\sum_{\begin{subarray}{c}\varepsilon_{k}\leq\varepsilon_{k+1},\\ k=1,...,i-1\end{subarray}}\sum_{\begin{subarray}{c}\varepsilon_{l}\geq\varepsilon_{l+1}\geq\varepsilon_{0},\\ l=i+1,...,N-1\end{subarray}}[P_{i-1}(\tilde{e}_{i-1}^{\varepsilon_{i-1}},\cdots,\tilde{e}_{1}^{\varepsilon_{1}}),P_{N-i+1}(\tilde{e}_{i+1}^{\varepsilon_{i+1}},\cdots,\tilde{e}_{0}^{\varepsilon_{0}})]_{q},$
where $(\varepsilon_{i-1},\varepsilon_{i+1})\in\{(+,-),(-,+),(+,+)\}$. Let us denote each summand on the RHS of (9.64), corresponding to a sequence $\vec{\varepsilon}=(\varepsilon_{i-1},\cdots\varepsilon_{1},\varepsilon_{i+1},\cdots\varepsilon_{N},\varepsilon_{0})$, by $P_{\vec{\varepsilon}}$.
It follows from Lemma 9.11 and (9.37) that
(9.65) $P_{i-1}(\tilde{e}_{i-1}^{\varepsilon_{i-1}},\tilde{e}_{i-2}^{\varepsilon_{i-2}},\cdots,\tilde{e}_{1}^{\varepsilon_{1}})=$ $\ [\tilde{e}_{i-1}^{\varepsilon_{i-1}},P_{i-2}(\tilde{e}_{i-2}^{\varepsilon_{i-2}},\cdots,\tilde{e}_{1}^{\varepsilon_{1}})]_{q},$
(9.66) $P_{N-i+1}(\tilde{e}_{i+1}^{\varepsilon_{i+1}},\tilde{e}_{i+2}^{\varepsilon_{i+2}},\cdots,\tilde{e}_{0}^{\varepsilon_{0}})=$ $\ [\tilde{e}_{i+1}^{\varepsilon_{i+1}},P_{N-i}(\tilde{e}_{i+2}^{\varepsilon_{i+2}},\cdots,\tilde{e}_{0}^{\varepsilon_{0}})]_{q}.$
Let us abbreviate $Y_{1}=P_{i-2}(\tilde{e}_{i-2}^{\varepsilon_{i-2}},\cdots,\tilde{e}_{1}^{\varepsilon_{1}})$ and $Y_{2}=P_{N-i}(\tilde{e}_{i+2}^{\varepsilon_{i+2}},\cdots,\tilde{e}_{0}^{\varepsilon_{0}})$. Then
(9.67) $P_{\vec{\varepsilon}}=$ $\ [[\tilde{e}_{i-1}^{\varepsilon_{i-1}},Y_{1}]_{q},[\tilde{e}_{i+1}^{\varepsilon_{i+1}},Y_{2}]_{q}]_{q}$
(9.68) $=$ $\ (1-q)[\tilde{e}_{i-1}^{\varepsilon_{i-1}},[\tilde{e}_{i+1}^{\varepsilon_{i+1}},Y_{1}Y_{2}]_{q}]_{q}.$
If $(\varepsilon_{i-1},\varepsilon_{i+1})=(+,+)$, then (9.43) in Lemma 9.12 now implies that $[P_{\vec{\varepsilon}},F_{i}]_{q^{-2}}=0$. Otherwise, $[[P_{\vec{\varepsilon}},F_{i}]_{q^{-2}},x^{-}_{i,r}]=0$ follows directly from (9.44)–(9.45) in Lemma 9.13. ∎
9.7. Factorization theorem in higher rank
We can now generalize the factorization and coproduct theorems to higher ranks.
Corollary 9.15.
Let ${\mathfrak{g}}$ be of type $\mathsf{A}$. Then:
(9.69) $\eta_{\mathbf{s}}(\boldsymbol{\grave{\Theta}}_{i}(z))\equiv$ $\ \xi_{\mathbf{s}}(\boldsymbol{\grave{\Theta}}_{i}(z))\boldsymbol{\phi}_{i}^{-}(z^{-1})\boldsymbol{\phi}_{i}^{+}(Cz){\mod U_{+}}[\negthinspace[z]\negthinspace],$
(9.70) $\Delta_{\mathbf{s}}(\boldsymbol{\grave{\Theta}}_{i}(z))\equiv$ $\ \eta_{\mathbf{s}}(\boldsymbol{\grave{\Theta}}_{i}(z))\otimes\eta(\boldsymbol{\grave{\Theta}}_{i}(z))\quad\mod{(U_{-}\otimes U_{+}+U_{+}\otimes U_{-})}[\negthinspace[z]\negthinspace].$
Proof.
If $\mathbf{s}=0$, the corollary follows from Corollary 4.11 and Theorem 9.10. Let us explain this more precisely. In analogy to (3.4), the defining relations of $\mathcal{O}_{q}(\widehat{{\mathfrak{g}}})$ imply that
(9.71) $\boldsymbol{\grave{\Theta}}_{i}(z)=1+\frac{c_{i}^{-1}Cz^{2}\left(z^{-1}[A_{i,-1},\boldsymbol{A}_{i,+}(z)]_{q^{-2}}-q^{-2}[A_{i,0},\boldsymbol{A}_{i,+}(z)]_{q^{2}}\right)}{1-Cz^{2}}.$
By Proposition 9.10, we can write $\eta(\boldsymbol{A}_{i,+}(z))=\iota_{i}\eta(\boldsymbol{A}_{+}(z))+R(z)$ with $R(z)\in{U_{d_{i}\geq 1,+}}[\negthinspace[z]\negthinspace]$. Moreover, $\eta(A_{i,-1})=\iota_{i}\eta(A_{-1})+R_{-1}$, with $R_{-1}\in U_{d_{i}\geq 1,+}$, and $\eta(A_{i,0})=\iota_{i}\eta(A_{0})$. We call $R(z)$ and $R_{-1}$ “remainder terms”. Since the elements $\iota_{i}\eta(A_{-1})$ and $\iota_{i}\eta(A_{0})$ contain precisely homogeneous terms of degree ±$\pm\alpha_{i}$, multiplying them with remainder terms yields an element of $U_{+}$. Similarly, a product of remainder terms lies in $U_{+}$. It follows that
(9.72) $\eta(\boldsymbol{\grave{\Theta}}_{i}(z))\equiv$ $\ \iota_{i}\eta\left(1+\frac{c_{i}^{-1}Cz^{2}\left(z^{-1}[A_{-1},\boldsymbol{A}_{+}(z)]_{q^{-2}}-q^{-2}[A_{0},\boldsymbol{A}_{+}(z)]_{q^{2}}\right)}{1-Cz^{2}}\right)$
(9.73) $=$ $\ \iota_{i}\eta(\boldsymbol{\grave{\Theta}}(z))\equiv\iota_{i}(\boldsymbol{\phi}^{-}(z^{-1})\boldsymbol{\phi}^{+}(Cz))=\boldsymbol{\phi}_{i}^{-}(z^{-1})\boldsymbol{\phi}_{i}^{+}(Cz)$
modulo ${U_{+}}[\negthinspace[z]\negthinspace]$. In the third equivalence, Corollary 4.11 was used.
The case of non-standard parameters $\mathbf{s}\neq 0$ follows from a higher rank analogue of (2.24) and Theorem 7.5, using the fact that $\overline{\Delta}_{\mathbf{s}}(U_{+})\subseteq{\mathbb{C}}\otimes U_{+}$. The group-like property of the coproduct is deduced in the same way as in the proof of Theorem 7.5. ∎
Thanks to Corollary 9.15, Corollaries 5.1 and 5.5 carry over to higher rank, as formulated in the introduction (Theorem 1.3).